A Complete Bibliography of Philosophy of Science: 2010–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

27 January 2020 Version 1.23

Title word cross-reference

$110.00 [756]. $35.00 [667]. $74.00 [640]. $78.00 [641]. _{ST} [142]. F [145].

-Statistics [145].

2016 [743].

83 [643]. 84 [632].

= [556].

Abductive [544]. Ability [712, 716]. Absolute [536, 585]. Abstract [271]. Abstraction [493, 525, 687, 189, 101]. Acceleration [521]. Acceptance [230]. Accepting [484]. Access [165]. according [357]. Account [664, 151, 132, 736, 235, 90, 213, 311, 567]. Accounts [665, 198, 784].

Accuracy [543, 180, 113, 129, 812, 640]. Accuracy-Firsters [543]. Achinstein [309]. Acidity [150]. Acquire [245]. Across [804, 740]. Act [775, 693]. Ad [105]. Adaptation [66, 528]. Adaptations [532, 99]. Addiction [421]. Address [743]. Addressing [679]. Adequacy [347]. Adequate [201]. Adjusting [723]. Adolf [306]. Adrian [115, 359]. Advancement [756]. Adventures [264]. Advisor [167]. Aesthetics [726]. Afraid [178]. Against [566, 605, 428, 364, 597, 44]. Agents [416]. Aggregating [298]. Aggregation [713]. Aggression [307]. Agree [777]. Agreeing [222]. Agreement [311]. Ahistorical [787]. Aims [741, 775]. Alcino [297]. Alex [382]. Algebra [628]. Algebraic [798, 170]. Algorithms [215]. Allan [395]. Allhoff [10]. Alspector [10]. Alspector-Kelly [10]. Alternative [32, 484, 238]. Alternatives [408]. Altruism [694]. Ambiguity [355, 290]. Analogical [721]. Analogs [693]. Analogy [70, 26, 69]. Analyses [172, 35]. Analysis [627, 470, 517, 166]. Angela [741]. Animal [692, 424]. Animals [380]. Answers [715]. Anthology [10]. Anthony [297, 370]. Anthropomorphism [610]. Anticommons [161]. Antilikelihood [535]. Antirealism [337, 73]. Antirealist [767]. Antireductionist [651]. Antiscience [646]. Antony [114]. Any [67]. Anything [270]. Applicability [198]. Application [737]. Applications [65, 139]. Applying [187]. Appraising [233]. Apprentices [264]. Approach [32, 514, 234, 614, 471, 798, 511, 750, 69, 432, 722]. Approaches [495]. Approximation [117]. Archaeology [577]. Argon [21]. Argument [711, 570, 113, 4, 482, 728, 651, 3, 69]. Arguments [286, 413, 617, 585]. Arnold [295]. Arntzenius [273]. Arrow [652, 255, 279]. Arrows [386]. Art [726]. Artifact [419]. Artificial [737]. Artificiality [328]. Ascent [704]. Aspect [478]. Aspects [671, 412]. Assertion [539]. Assessments [561]. Assumption [772, 423]. Assumptions [350]. Astronomy [383, 805]. Astrophysical [676]. Asymmetry [350, 558]. Atoms [801]. Attachment [226, 733]. Attitudes [229]. Attribution [334]. Augmented [449]. Author [163]. Autonomous [51, 762]. Autonomy [474]. Avoidance [733]. Awareness [795]. Awkward [77]. Axiomatic [639].

Bacciagaluppi [114]. Background [286, 490]. Bad [448]. Bain [655]. Bar [454]. Bar-Hillel [454]. Bar-Hillel/Carnap [454]. Bardon [359]. Barking [23]. Barrett [115]. Base [31]. Base-Rate [31]. Based [184, 511, 422, 317]. Baseline [660]. Basic [320]. Basis [371, 159, 811, 426]. Bateson [152]. Batterman [366]. Bayes [627]. Bayesian [204, 626, 467, 535, 13, 387, 224, 266, 748, 267]. Bayesianism [664, 14, 15]. Be [391, 448, 151, 767, 460, 116, 501, 152, 265, 74, 178, 67]. Beauty [468]. Behaving [813, 641]. Behavior [307, 434, 317]. Behavioral [512, 611]. Being [149]. Belief [225, 139, 752, 751]. Beliefs [425]. Bending [552]. Best [32, 749]. Best-Alternative [32]. Better [391, 174, 728, 571]. Between [61, 362, 323, 77, 738, 458, 435, 635, 87, 173, 524]. Beyond [196, 235, 198, 473, 231, 583, 94]. Bias [610, 680, 442, 250]. Biased [416].

Bickle [297]. Big [362]. Biography [296]. Biological [659, 588, 39, 513, 785, 427, 708, 319, 691, 152, 242, 431, 274, 591, 516, 474]. Biologically [564]. Biologists [686, 240]. Biology [146, 654, 656, 526, 690, 786, 688, 687, 206, 746, 700, 119, 362, 38, 455, 315, 809, 396, 729]. Biomedical [151, 163]. Birthday [306]. Black [220, 485, 485, 702]. Blackmore [9]. Blurring [410]. Body [22]. Bohm [258, 434]. Boltzmann [753]. Boltzmannian [209, 438, 621]. Bond [23]. Bonding [801]. Book [540, 667, 488, 357, 273, 359, 383, 718, 371, 184, 370, 122, 308, 382, 114, 130, 183, 307, 296, 384, 294, 104, 309, 210, 717, 578, 295, 139, 271, 367, 297, 11, 154, 369, 115, 413, 138, 640, 358, 9, 103, 193, 140, 396, 283, 395, 281, 368, 455, 282, 394, 456, 10, 641, 729, 730, 272, 541, 756]. Bottom [344]. Bottom-Up [344]. Boumans [578]. Boundedly [227]. Bounds [627]. Boxes [220, 209]. Boyer [717]. Boyer-Kassem [717]. Brain [325, 271, 332, 420, 495, 491]. Brains [279]. Branching [253]. Brandon [87]. Break [609]. Breaking [168, 483]. Brief [359]. Broadbent [382]. Broglie [434]. Broken [81]. Brunswick [756]. Building [206]. Button [127].

C [122, 756]. Camera [271]. Can [331, 32, 448, 182, 19, 617, 67]. Cancer [659, 686]. Cannot [182]. Canonical [1]. Captures [271]. Cardinality [188]. Care [125, 813, 641]. Careful [614]. Carl [718]. Carnap [20, 454, 295]. Cartan [451]. Case [650, 147, 526, 236, 779, 501, 772, 148, 335, 360, 21, 223]. Cases [146, 334]. Casual [334]. Catastrophism [409]. Categorical [710]. Categories [321]. Causal [218, 532, 298, 147, 219, 698, 760, 514, 34, 385, 171, 287, 513, 714, 30, 772, 187, 90, 386, 761, 86, 608, 497, 658, 444, 368, 590, 815, 516, 564, 311, 217]. Causality [460, 33, 456]. Causally [450, 762, 533]. Causation [142, 546, 249, 99, 574, 690, 811, 400, 612, 444, 12, 173, 311, 622, 371, 358]. Cause [96, 216]. Causes [25, 425, 559, 42]. Causing [453]. Caution [263]. Cautionary [681]. Cell [241]. Centric [656]. century [807, 238, 395]. Certainty [587]. Ceteris [34, 638]. Chain [69]. Chains [82]. Challenge [287, 677, 744, 479]. Challenges [65, 540]. Challenging [73]. Chance [389, 462, 560, 797]. Chances [388]. Change [561, 65, 71, 64, 152, 528]. Chaos [439, 227]. Characterization [256]. Characterizing [96]. Cheap [417]. Chemical [305, 801, 59]. Chemistry [56, 259, 58]. Chien [756]. Chirimuuta [487]. Choice [628, 772, 424, 79, 691, 377]. Christopher [308, 293]. Churchland [271]. Citation [344, 778]. Clade [547]. Clades [338]. Claim [78]. Claims [147, 128]. Classes [721]. Classic [541]. Classical [255, 434, 365, 620]. Classification [383, 780, 305, 805, 625]. Clauses [34]. Clerk [309]. Climate [561, 236, 529, 63, 165, 65, 66, 71, 64, 530, 528]. Clines [338]. Clinical [352, 446]. Closed [96]. cloth [640, 641, 756]. Clusters [632, 338, 593]. Coarse [714, 437]. Coarse-Grained [714]. Coarse-Graining [437]. Coevolution [303]. Cognition [692, 23]. Cognitive [378, 610, 292, 228, 229, 611, 24, 548, 673, 172, 501, 134, 101, 55]. Cognitive/Evolutionary [548]. Coherence [418]. Coin [728].

Coin-Tossing [728]. Coincidences [93]. Collaboration [391, 717, 163, 472, 360, 647]. Collaborators [777]. Collection [630]. Collective [717]. Color [481, 487]. Commensuration [442]. Comment [631, 484]. Comments [752]. Commercialization [162]. Common [96, 470]. Communication [496]. Communism [573]. Communities [712, 653]. Community [428]. Comparative [70, 469, 71]. Comparativism [616]. Compatibilist [664]. Compelling [263]. Competition [583]. Completeness [401, 802]. Complex [344, 504, 503, 505, 513, 343, 345, 72]. Complexity [415, 103, 582]. Comprehensive [421]. Computation [261, 54, 567]. Computational [27]. Computational-Level [27]. Computationalism [53]. Compute [735]. Computer [398]. Con [9]. Concept [314, 700, 608, 658]. Conception [674, 801, 727]. Conceptions [507]. Concepts [320, 534, 151, 245, 187, 314]. Conceptual [540, 370, 806]. Concerning [256]. Condition [179, 342, 559]. Conditionalization [180]. Conditionals [774]. Conditioning [250]. Conditions [99, 460, 472]. Conference [114]. Confidence [561]. Confirmation [275, 551, 535, 63, 62, 268, 536, 95, 411, 267]. Confirmational [222]. Conflations [90]. Confrontation [186]. Confronting [512]. Confusions [28]. Conglomerability [180]. Conjunction [2]. Connection [655]. Connections [620]. Conor [717]. Consciousness [248, 795]. Consensus [805, 722, 383]. Consent [443]. Consequences [15]. Conservation [93, 700, 237, 620]. Considerations [469]. Consilience [171, 414]. Consistency [518, 517]. Constancy [481, 329]. Constitution [544]. Constraining [170]. Constraints [120, 659, 699, 93, 242, 492]. Construct [527, 796, 491]. Constructed [662]. Construction [7, 691, 394, 808]. Constructions [794]. Constructive [550]. Contemporary [163]. Content [576]. Contextualism [76]. Continental [502]. Contingency [428, 606, 480]. Contingent [599]. Contingentist [239]. Contingentist/Inevitabilist [239]. Contractarian [221]. Contradictions [162]. Contrastive [675, 734]. Control [659, 720]. Controversies [165, 612, 756]. Controversy [383, 805]. Conventionalism [243, 337]. Conventionality [277]. Convergence [387]. Convergent [186]. Cooperation [417, 109]. Cooperative [290]. Coordination [795]. Corporate [680]. Correct [289]. Correlation [711, 95]. Cosmic [410, 28]. Cosmology [617, 193]. Could [116, 265]. Counterexamples [414]. Counterfactual [443, 558, 497]. Counterfactuals [124, 213]. Counternomic [594]. Counts [404]. Coupling [472]. CPT [655]. Crash [609]. Creation [324]. Credence [389, 812, 640]. Credences [468, 626, 300]. Criteria [679]. Criterion [737, 157]. Critical [526, 270, 293]. Criticism [556]. Critique [124, 182, 164, 280]. Cross [574]. Cross-Unit [574]. Crossroads [114]. Cultural [540]. Culture [418, 534]. Cure [170]. Curie [519, 520, 208]. Curves [638]. Cycles [514, 504, 503, 505].

D [103]. Daniel [384, 729]. Dark [706]. Darwinian [361, 140, 110]. Daston [541]. Data [328, 22, 404, 670, 49, 165, 671, 193, 405, 458, 412, 362, 21, 649, 48, 630, 656]. Data-Centric [656]. Data-Intensive [412]. Davey [209]. David [115, 603, 138]. Deal [286]. Dear [104]. Debate [646, 196, 256, 350, 430, 338, 239, 624, 409]. Debates [73]. December [632]. Deception [322]. Decision [561, 79, 456]. Decisions [205]. Decomposition [242]. Deep [243]. Defeasibility [678]. Defense [666, 199, 796, 311]. Definable [710]. Defining [361]. Definition [136, 796]. Degrees [239]. Deidealization [745]. Demand [638, 165, 201]. Demarcation [673]. Demise [428]. Democratic [443, 601, 272]. Democratizing [681]. Demon [264, 82]. Demystifying [169]. Dependence [113, 171, 558, 137]. Dependency [312, 690]. Depth [714, 16]. Deriving [433]. Destroyed [67]. Detecting [146]. Detrimental [599]. Development [629]. Developmental [146, 503, 78, 72]. Device [245]. Diagnostic [423]. Diagrams [240]. Dialogues [683]. Dick [383]. Did [797]. Diego [54]. Difference [599, 117, 409]. Difference-Makers [599]. Differences [501, 740]. Difficulty [247]. Digital [685]. Dimension [101]. Dimensional [227]. Direct [91, 747]. Disagreement [287, 87]. Disagreements [738]. Discernibility [435]. Discerning [177]. Disclosure [390]. Discontents [508]. Discordant [352]. Discoveries [804]. Discovering [386]. Discovery [353, 805, 108, 148, 21, 383]. Discrimination [647, 423]. Discussing [683]. Discussion [102, 214, 270, 26]. Disease [789, 333]. Disjunctive [650]. Dissatisfaction [68]. Dissent [599]. Dissociations [134]. Dissolve [617]. Dissolving [602]. Distinguishing [153]. Distortion [793]. Distribution [1]. Diversity [712, 461, 716, 530]. Divide [109]. Division [378, 292, 673, 379]. DNA [108]. Do [661, 391, 777, 124, 634, 245, 66, 257, 685, 686, 255, 240, 279, 502]. Does [82, 609, 529, 728, 761, 201, 407, 263, 100, 58]. Dog [23]. Doing [332]. Dominated [449]. Dominique [756]. Douglas [371]. Down [546, 609]. Downward [249, 690]. Drift [185]. Drugs [680]. Dualities [524]. Duncan [370]. Dupré [729]. Dutch [413]. Dwindling [268]. Dynamic [470]. Dynamical [256, 798, 351, 589, 432, 516, 88]. Dynamics [546, 363, 13, 738, 495].

Easy [745]. Ecological [699, 653]. Ecology [698, 316, 428]. Economics [459, 392, 472, 701, 578]. Ecphory [696]. eds [717, 9, 10, 729]. Eduardo [632]. Effect [353, 205, 733, 529]. Effective [472]. Effectiveness [174]. Effects [765, 562]. Efficiency [788, 192, 135, 701]. Eikonic [674]. Einstein [68]. Election [441]. Electroweak [737]. Elements [289]. Eliminative [85, 362]. Elliott [130, 730, 488, 484]. Elsewhere [353]. Elusive [159]. Embedded [429]. Embodiment [421]. Emergence [417, 718, 705, 749, 317]. Emergent [357, 112]. Emery [754]. Emotional [693]. Emotions [794, 662]. Empathy [270, 473]. Empirical [254, 430, 347, 49, 165, 671, 452, 87, 157]. Empiricism [20]. Enabling [599]. End [261]. Engineering [609, 297].

Engines [377]. Engram [696]. Enriched [648]. Ensembles [64]. Entanglement [67]. Entropy [136, 59]. Environment [563, 314]. Environmental [384, 99, 700, 166, 814, 283]. Environments [99]. Epidemiological [382]. Epidemiology [803, 382]. Epiphenomenalism [790]. Epistemic [378, 739, 646, 780, 415, 712, 769, 673, 779, 569, 463, 101, 299, 545, 408, 3, 379]. Epistemically [599, 484]. Epistemological [671]. Epistemology [7, 743, 217]. Equations [34, 612]. Equilibrium [753, 471, 227, 438, 621]. Equipped [434]. Equivalence [521, 191, 710, 458, 596]. Ernst [9]. Erosion [677]. Erratum [84, 211, 372, 457, 632, 643]. Error [665]. Eschatological [392]. Essay [306, 566, 485]. Essays [717]. Essences [37]. Essential [166]. Essentialism [305, 39, 30, 38]. essentialist [244]. Estimating [145]. Ethics [700]. Ethnic [245]. Etiology [318]. Evaluate [680]. Evaluating [672]. Evaluation [538, 582]. Evasion [327]. Everett [357, 115]. Everyday [150]. Everyone [43, 799]. Everything [475, 729]. Evidence [648, 448, 184, 179, 147, 384, 570, 767, 672, 393, 352, 629, 193, 405, 374, 444, 95, 376, 814, 456, 557, 722, 649, 586, 414, 309]. Evidence-Based [184]. Evidential [461, 537, 577]. Evolution [142, 540, 320, 41, 256, 510, 563, 25, 511, 506, 79, 224, 694, 23, 322, 449, 110, 183]. Evolutionary [120, 243, 534, 504, 375, 688, 548, 606, 426, 26, 301, 42, 782, 72]. Evolving [321, 109]. Exaggerated [428]. Exchange [205]. Exclusivity [583]. Exercise [348]. Expanded [506]. Expect [218]. Expected [180]. Experience [329, 330, 421]. Experiences [328]. Experiencing [755]. Experiment [424, 217]. Experimental [219, 328, 736, 740, 720, 315, 721]. Experimentation [598]. Experimenter [221]. Experiments [565, 131, 329, 683, 299, 395, 362, 807]. Expert [528]. Expertise [713, 712]. Explain [182, 686, 58]. Explaining [674, 478, 343]. Explanation [674, 689, 588, 551, 132, 526, 765, 636, 4, 375, 342, 287, 106, 400, 118, 675, 5, 496, 346, 497, 556, 351, 213, 431, 590, 763, 591, 60, 734, 88]. Explanations [401, 288, 650, 398, 382, 688, 714, 93, 185, 366, 345, 497, 213, 98]. Explanatoriness [214]. Explanatory [774, 664, 153, 354, 650, 494, 126, 214, 287, 687, 714, 399, 675, 195, 589, 80, 242, 649, 227, 16]. Explicating [546]. Exploratory [764, 362]. Exposed [771]. Expressions [693]. Extended [23]. Extension [306]. Extensions [339]. Extensive [406]. External [445]. Extrapolation [70, 69, 721].

F [771, 641, 142]. Face [79]. Factors [532, 599, 333]. Failure [208]. Faithfulness [218, 772]. Fallacy [31, 591]. Favor [724]. Featuring [514]. Feel [121]. Feigl [83]. Fictionalist [746]. Fictions [153, 154]. Field [565, 370, 806, 168, 550, 578, 437, 68, 67]. Fields [190, 798]. Fifty [541]. Files [634]. Financial [343]. Finding [192]. Fine [570]. Fine-Tuning [570]. Fingerprints [71]. Finite [483]. First [328, 375, 496, 72]. First-Person [328]. Firsters [543]. Fish [529]. Fit [592]. Fitelson [137]. Fitness [507, 128, 363, 505, 708, 509, 86, 426, 624, 12]. Flash [278]. Flatness [617]. Flows [729]. Focused [95]. Fodor [33]. Following [320]. Forbidden [501]. Force [207]. Forces [42]. Foregrounding [490]. Forest [529]. Forge [11]. Forging [155]. Formal [518, 269]. Formulated [78, 58]. Formulation [703]. Forster [414]. Forward [196]. Foundation [709]. Foundations [308, 566, 138, 394, 808]. Founded [151]. Framework [370, 609, 806, 269, 404, 625, 263, 238]. Framing [728]. Frank [273, 356]. Franklin [395, 565, 108]. Free [448, 584, 479]. Frequentist [31]. Friedman [75, 556]. Fritz [10]. Frost [295]. Frost-Arnold [295]. Fruits [109]. Function [302, 562, 787, 363, 232, 789, 175, 720, 792]. Functional [788, 325, 194, 135, 420, 422, 35, 246, 750, 311]. Functionality [215]. Functions [302, 24, 605, 735]. Fundamental [371, 811]. Fundamentality [652, 341]. Funded [679, 680, 681]. Fungi [304]. Future [65, 144].

G [308]. Galaxy [706]. Galileo [683, 452, 684]. Game [603]. Games [390, 470]. Gap [665, 621]. Garber [393, 571]. Garber-Style [393]. Gases [209]. Gauge [613, 737, 452, 635, 522, 524, 643]. Gendered [778]. Gene [324]. General [433, 614, 138, 252]. Generality [698]. Generalized [562, 253, 575]. Generic [335]. Genes [363]. Genetic [215, 185, 813, 641]. Genetics [294]. Genic [87]. Genomics [782]. Genotypic [87]. Gentle [280]. Geodesic [252]. Geometrized [340]. Geometry [798, 432, 277, 566]. Germs [363]. Get [151, 799]. Getting [158]. Gibbs [136, 1, 753]. Gibbsian [621]. Gillett [718]. Global [66]. GMO [646]. Go [625, 261]. Goals [262]. Gödel [523]. Godfrey [140, 396]. Godfrey-Smith [140, 396]. Going [386]. Good [130, 355]. Got [33]. Gould [349]. Graded [543]. Grades [285]. Grained [714]. Graining [437]. Grammar [535]. Graphs [69, 516]. Gravitation [637, 138, 340]. Gravitational [521]. Gravity [616, 193]. Gray [296]. Greatly [428]. Greg [295]. Griffiths [294]. Ground [124, 206]. Grounding [157]. Groundings [101]. Grounds [593, 632, 263]. Group [636, 437, 345, 346, 704]. Groups [243, 574]. Grünbaum [306]. GRW [278]. Gualtiero [567]. Guide [667, 565, 560, 569, 358]. Guides [551]. Guido [114]. Guilt [511].

H [122]. Hacking [695, 586]. Hall [358]. Halvorson [280]. Hamilton [507, 509]. Handicap [419]. Handle [125]. Hans [280]. Happen [683]. Haptic [498]. Hard [772]. Harm [778]. Harper [193]. Harvard [295]. Hawkmoth [529]. Heads [391]. Health [135, 381]. Helen [122]. Helps [174]. Henri [296]. Heredity [128, 152]. Heritability [602, 533]. Heterogeneous [420]. Heterokaryotic [304]. Hidden [386]. Hierarchical [13, 111]. Hierarchies [515, 415]. Higgs [353, 170]. Higher [40, 762]. Hillel/Carnap [454]. Historical [152, 768, 10, 145, 756]. Historically [767]. Historicity [312]. History [359, 548, 480]. HIV [761]. Hoc [105]. Hodgkin [612]. Hole [482]. Holes [339, 463, 485]. Holism [29, 430, 677]. Homeostasis [40]. Homology [608]. Honor [306]. How-Possibly [398].

Howick [184]. Hoyningen [281]. Hoyningen-Huene [281]. Huene [281]. Human [692, 183, 307, 495, 244]. Humans [502]. Hume [45, 32]. Humeanism [749]. Humeans [799]. Humility [467]. Hunt [449]. Huxley [612]. Huygens [464, 451]. Hypotheses [105]. Hypothesis [254, 436, 413, 424, 583, 223, 706]. Hypothetical [195].

Ian [695]. IBE [399]. Iconic [576]. I’d [74]. Idea [609]. Ideal [479]. Ideality [523]. Idealization [400, 117, 195, 684, 741]. Idealizations [76]. Identity [190, 574, 350, 8]. If [202, 270, 74]. II [15]. Illumination [481]. Illusion [330]. Image [389]. Images [17]. Imaging [422]. Imperfect [236]. Implications [142]. Import [537]. Importance [659]. Imprecise [225, 448]. Inaccuracy [14, 15]. Inaugurating [106]. Incentive [573]. Incest [733]. Inclusive [708, 509, 624]. Incoherence [543]. Incompatibility [516]. Increasing [722]. Increasingly [128]. Independence [678, 179]. Independent [557]. Indeterminate [300]. Indexically [653]. Indigenous [779]. Indirect [91]. Indiscernibles [8]. Indispensability [589]. Indistinguishability [219]. Indistinguishable [177]. Individual [317]. Individual-Based [317]. Individualism [316, 430]. Individuality [505, 427, 606, 607, 784]. Individuals [304, 350]. Induction [43, 572, 45, 46, 747, 773, 44, 584]. Inductions [768]. Inductive [231, 631, 639, 584, 285, 3, 230, 479]. Industrial [598]. Industry [679, 680, 681]. Industry-Funded [679, 680, 681]. Inertial [521, 464, 684]. Inevitabilist [239]. Inextendibility [615]. Inference [85, 4, 772, 747, 639, 748, 217]. Inferences [362]. Inferential [159, 723]. Inferior [484]. Inferring [303, 368, 815]. Infinite [651]. Infinitesimal [746]. Information [161, 448, 390, 324, 308, 454, 363, 286, 440, 90, 691, 263, 274]. Informational [666]. Informed [443, 722]. Inherent [582]. Inheritance [152]. Inherited [78]. Injustice [780, 779]. Inner [247]. Inquiry [776, 11]. Ins [315]. Insights [659, 761]. Integrative [784, 492]. Intellectual [779, 161]. Intelligibility [104]. Intemann [631]. Intended [680]. Intensive [412]. Intentional [187]. Intentionality [392]. Interaction [421]. Interactions [173]. Interactive [801, 99, 559]. Interdependence [699, 698, 700]. Interdisciplinary [472]. Interest [470]. Interests [225]. Interpretation [604, 402, 618, 57, 357]. Interpretations [614]. Interpreting [450, 533, 210]. Interpretive [748]. Intersectionality [450]. Intertheoretic [744]. Intervention [696]. Interventionist [526, 622]. Interventions [594]. Intimate [620]. Intransigently [416]. Intrinsic [37, 175]. Intrinsicalism [6]. Introduction [730, 294]. Introspection [327, 247]. Intuitive [311]. Invariance [214, 77, 618, 655]. Investigate [307]. Irrelevant [2]. Isaac [309, 193]. Issues [430]. Itagaki [9].

J [667, 632, 383, 297, 729, 541]. Jacob [270]. James [309, 643, 790, 23]. Jeffreys [276, 178]. Jeremy [184, 296]. Jigsaw [95]. Jim [752]. John [297, 729, 11, 9]. Joint [132]. Jonathan [115, 655]. Joyce [113, 137].

Judgment [528, 311]. Judgments [298, 167]. Justice [650, 552]. Justification [32, 392, 14, 15, 209, 773, 148]. Justificatory [234]. Justifies [399].

Kant [808, 291, 394]. Kantian [75]. Karen [666]. Karola [294]. Kassem [717]. Kellert [122]. Kelly [10, 192]. Kenneth [122, 641]. Kent [115]. Kevin [130, 730]. Kills [500]. Kin [508]. Kind [625, 564]. Kinda [355]. Kinds [678, 764, 73]. Kitcher [272]. Knew [74]. Know [232, 66]. Knowability [301]. Knowing [17]. Knowledge [717, 501, 335, 553]. Known [330]. Knox [451]. Kuhn [30, 26, 541]. Kuhnian [164]. Kutach [371].

L. [358]. Labor [378, 292, 673, 379]. Laboratory [578]. Ladenness [412]. Landreth [297]. Landscape [271, 420]. Landscapes [378, 688]. Lange [124]. Language [29, 113, 137]. Laplace [264]. Latent [721]. Laura [210]. Law [375]. Laws [92, 124, 139, 93, 19, 640, 812, 749, 33, 237, 620, 349, 752, 751]. Lawton [428]. Laypersons [673]. Learn [331]. Learning [417, 403, 21, 227]. Led [521]. Left [586]. Legitimacy [311]. Leitgeb [129]. Leonelli [656]. Leopold [700]. Less [545]. Lessons [352, 688, 215]. Letting [625]. Level [27, 60]. Levels [249, 762, 143, 715, 5, 590]. Lewens [540]. Lewis [667, 603]. License [407]. Life [804, 504, 503, 736, 505, 670, 720, 455]. Light [782]. like [121]. Likelihood [535, 286, 414]. Limit [651]. Limitations [111]. Limits [32, 513, 651]. Lindley [178, 276, 223]. Lineage [143, 783]. Lineages [143]. Liouville [82]. Lisa [756]. Literal [614]. Little [130]. Local [67]. Localization [175, 242]. Logarithmic [724]. Logic [126, 80, 702]. Logics [726]. Longino [122]. Look [126, 353]. Looking [234]. Loops [769]. Loose [500]. Lorraine [541]. Love [74]. Lunch [584].

M [271]. Mach [9]. Macrobehavior [346]. Macroeconomics [392]. Magnetic [422]. Maintenance [418]. Make [802]. Makers [599]. Makes [104, 409, 167]. Making [547, 592, 553]. Malament [138]. Manage [680]. Manifold [253]. Manipulability [214, 444]. Manual [488]. Many [148, 240, 115]. Mapping [198]. Mappings [688]. Marc [10]. Marcel [578]. Margaret [486]. Mark [666]. Markets [343]. Markov [559]. Marr [27]. Martín [631]. Martinez [632]. Mass [521, 616]. Masthead [466, 476, 531, 542, 555, 568, 579, 580, 623, 633, 644, 657, 668, 707, 719, 732, 742, 757, 758]. Matching [481]. Material [235, 747, 768]. Mathematical [591]. Mathematics [486, 62, 198, 156, 293]. Matter [278, 239, 761, 425, 455, 58]. Matters [218, 332, 424, 117, 564, 311]. Matthewson [234]. Maudlin [282]. Maximally [722]. Maxwell [82, 637, 451, 309]. Mayo [717]. Mayo-Wilson [717]. Mazviita [487]. McGrew [10]. McKaughan [484]. Meanings [538]. Measure [197, 708, 517, 209]. Measurement [402, 578, 486, 769, 220, 586, 157]. Measurements [222]. Measures [538, 354, 585, 411, 267]. Measuring [385, 14]. Mechanical [207].

Mechanics [136, 168, 8, 753, 68, 365, 209, 620, 667, 308]. Mechanism [689, 604, 194, 102, 170, 763, 238]. Mechanisms [92, 659, 804, 269, 514, 19, 189, 429, 422, 444, 516, 369]. Mechanistic [120, 401, 493, 515, 132, 526, 762, 235, 400, 567, 431]. Medical [381, 250, 553, 681]. Medicine [334, 789, 184]. Medium [17]. Meet [768]. Meets [79]. Melo [631]. Memes [363]. Memory [696]. Mendeleev [802]. Mental [666]. Mereological [57]. Message [17]. Meta [672, 584]. Meta-induction [584]. Meta-Research [672]. Metainductive [773]. Metaphysical [394, 808]. Metaphysically [237]. Metaphysics [667, 478, 387, 364, 407, 311]. Method [500, 309, 193]. Methodological [352, 316, 430, 677]. Methodologies [624]. Methodology [459]. Methods [327, 328, 415]. Metric [100]. Michael [717, 485, 183]. Micro [112]. Micro-Ontology [112]. Microbial [455]. Microfoundations [392]. Microscopes [695]. Microscopic [437]. Microstructure [305]. Mill [43]. Mind [678, 152, 491, 453, 621]. Mindreading [380]. Minds [424]. Minimal [288, 403, 366]. Minimizing [15]. Miracles [4]. Misleading [767]. Mismeasure [795]. Missing [602, 529]. Mitchell [103]. Mixed [765, 60]. Mixed-Effects [765]. Mixed-Level [60]. Mobile [89]. Modal [302]. Modality [785]. Model [288, 538, 41, 515, 292, 737, 366, 472, 746, 511, 155, 422, 582, 530, 702, 421, 94]. Model-Based [511, 422]. Model/World [155]. Modeling [676, 645, 660, 765, 514, 99, 513, 235, 206, 156, 315, 492, 157]. Models [344, 52, 494, 236, 403, 234, 233, 486, 63, 764, 405, 386, 66, 781, 64, 195, 351, 87, 317, 154]. Modest [626]. Modularity [196, 677, 709]. Molecular [689, 786, 687, 118]. Molecules [801]. Monism [73]. Monism/Pluralism [73]. Monist [497]. Monophyly [40]. Moral [425]. Morality [510]. Morrison [486]. Most [547, 447]. Motion [614]. Motivated [767]. Motivation [618]. Movements [779]. Moving [196]. Much [478, 529, 761]. Muller [418]. Multidimensional [143, 625]. Multigenerational [506]. Multilayer [495]. Multilevel [304, 143, 319]. Multilineal [506]. Multiple [262, 24, 775, 234, 785]. Multiple-Models [234]. Multiverse [357]. Mushroom [607]. Mutual [583]. Myopia [236]. Myth [251]. Myths [549].

Nagel [94]. Nagelian [94]. Nanay [182]. Narrative [763]. Nash [227]. Natural [678, 604, 102, 625, 140, 35, 73, 263, 394, 33, 291, 734, 808]. Naturalism [135, 74]. Naturalist [560, 74]. Nature [124, 104, 19, 244, 281, 394, 88, 808]. Neander [666]. Necessary [237]. Ned [358]. Need [391, 777, 245, 66]. Needs [673]. Negative [160]. Neo [75]. Neo-Kantian [75]. Neolithic [377]. Network [196, 494, 172, 495, 709]. Network-Oriented [172]. Networks [546, 344, 324, 363, 545]. Neural [270, 326, 709, 492, 791]. Neurobiology [55]. Neuroeconomics [275]. Neuroimaging [331]. Neuropsychology [134]. Neuroscience [609, 249, 24, 134, 351, 297]. Neutral [660, 688]. Newcomer [73]. Newton [309, 193, 43, 521, 207, 176, 587, 451]. Newtonian [138, 572, 41, 521, 616, 340]. Next [297]. Nicholson [729]. Nijhout [87]. Nina [754]. Ninetieth [306]. NJ [756]. No [249, 4, 403, 787, 380, 350, 745, 107, 46, 261, 584, 632, 643]. No-Go [261]. Noise [47, 48]. Nominalism [61]. Non [41]. Non-Newtonian [41]. Nonadditive [509]. Nonarbitrary [625]. Noncausal [346, 497, 213]. Nonepistemic [262, 539, 723]. Nonexplanatory [153]. Nonhuman [380]. Nonmodular [215]. Nonreductive [765]. Nonrelativistic [68]. Nonrepresentationally [233]. Norms [503, 162]. Norton [43, 45]. Novel [126, 348, 376]. Novelty [600]. Null [660, 424, 223]. Numerical [415].

Object [149, 634]. Objection [790, 601]. Objective [560, 14, 15]. Objectivity [679, 200, 789]. Objects [611, 257]. Obligations [408]. Observables [619]. Observation [200, 810, 148, 367]. Occam [786, 477]. Ockham [488, 192]. Off [586, 234]. Old [393, 629, 671, 556, 376]. Omissions [526]. One [391, 462, 478, 148, 107, 556, 227, 280]. One-Dimensional [227]. Only [74, 311]. Ontological [170]. Ontologies [77]. Ontology [667, 278, 172, 49, 112, 201]. Open [430]. Operation [689, 67]. Operationalizing [248]. Opinion [441]. Opinions [722]. Opposed [311]. Optimal [378, 79]. Optimality [584, 449, 591]. Optimizing [377]. Option [449]. Optogenetic [696]. Optogenetics [695, 697]. Ordinary [151]. Organic [259]. Organismic [426]. Organisms [323]. Organization [725, 189, 111]. Organizational [605]. Orgulity [204]. Oriented [172]. Ostwald [56]. Outcome [585]. Outcomes [120]. Overlooked [334]. Overrated [766]. Owen [643]. Oxford [667, 640, 641].

Packages [154]. Pain [270, 121, 473]. Palmarini [33]. paper [667]. Paradigm [361]. Paradox [2, 306, 181, 266, 276, 260, 223, 136, 454, 178]. Paribus} [34, 638]. Parity [90]. Partial [601]. Particle [77, 395, 807]. Partly [37]. Parts [257]. Path [312]. Pathway [658]. Pattern [195]. Patterns [51, 588, 49]. Paul [358, 294, 271, 281]. Payoffs [509]. Peculiar [702]. Peer [164, 442]. Peircean [399]. Perception [792]. Perceptual [795, 321, 487]. Performance [248]. Periodic [802]. Permutation [77]. Perpetuum [89]. Persistence [150]. Person [328]. Perspectival [645]. Perspective [304, 305, 196, 503, 13, 746, 408, 756, 455]. Perspectives [654]. Perspectivism [498]. Pessimism [704]. Peter [667, 104, 309, 140, 396]. Pettigrew [640, 129]. Phase [57, 744, 260]. Phenomena [611, 513, 461, 458, 50, 48]. Phenomenal [328, 329]. Philip [272]. Philosophical [656, 139, 11, 152, 740, 173]. Philosophy [18, 123, 203, 643, 359, 184, 803, 174, 183, 294, 809, 348, 810, 367, 207, 75, 9, 283, 685, 740, 455, 291, 814, 729, 384, 396, 282, 10, 718, 487, 284, 373, 465, 554, 642, 731, 382]. Phonetic [725]. Phylogenetics [429]. Phylogeny [143]. Physical [56, 566, 271, 798, 201, 567]. Physically [97]. Physics [76, 811, 437, 112, 395, 365, 282, 98, 807, 371]. Piatelli [33]. Piatelli-Palmarini [33]. Piccinini [567]. Pick [634]. Picking [586].

Piecewise [286]. Pills [447]. Pincock [293]. Placebos [447]. Plasticity [322]. Plato [271]. Platonism [61]. Plausible [202]. Pluralism [73, 784, 697, 122]. Pluralist [440]. Plurality [606]. Poincaré [296]. Point [638]. Polarization [538]. Policies [680]. Policy [174, 814, 167, 384, 103]. Political [147]. Polling [441]. Pollution [130]. Pool [773]. Population [314, 782, 317]. Populations [304, 313, 361, 140, 502]. Portable [171]. Positive [532]. Positivist [459]. Possibility [763]. Possible [387]. Possibly [398]. Post [244]. Post-essentialist [244]. Potential [612]. Potochnik [741]. Power [354, 494, 126, 160, 80, 649]. pp [667, 640, 641, 756]. Practical [274, 432]. Practice [7, 539, 427, 381, 743, 283]. Practices [778, 530]. Pragmatic [225, 469, 714, 458]. Pragmatist [374, 569]. Prairie [485]. Precaution [166]. Precautionary [384, 814]. Precise [223]. Predicting [174, 64]. Prediction [477]. Predictions [441, 802]. Preface [141, 759, 581, 669, 36, 397, 489, 212, 310]. Preferences [663, 459]. Preferential [226]. Preparative [131]. Presentism [254]. Presidential [743]. Press [667, 640, 641]. Primitive [201]. Principal [518, 478, 517]. Principia [176]. Principle [518, 519, 384, 478, 241, 521, 20, 419, 8, 517, 814, 684, 575, 208]. Principles [535]. Priority [72]. Privilege [299]. Pro [9]. Probabilism [113]. Probabilist [448]. Probabilistic [714, 326]. Probabilities [64]. Probability [142, 561, 96, 258, 469, 188, 368, 209, 815, 12, 587]. Problem [32, 595, 602, 197, 767, 482, 380, 393, 416, 614, 772, 445, 617, 795, 137, 152, 582, 330, 336, 376, 748]. Problems [249, 102]. Process [504, 70, 71]. Processes [689, 431, 301]. Processing [454]. Processual [729]. Producing [22]. Progress [750, 697, 423]. Projections [529]. Promising [32]. Proper [256]. Properly [406]. Properties [190, 58]. Property [632, 161, 593]. Prospects [170]. Prosser [755]. Prosthetic [52]. Proto [684]. Proto-Inertial [684]. Provide [263]. PSA [743]. Psychiatric [780, 423]. Psychological [329, 794, 694, 173]. Psychology [548, 796]. Psychometric [164]. Psychophysical [327]. Public [673]. Punishment [449]. Pursuit [776, 399]. Push [127]. Put [82]. Putting [571, 496]. Puzzle [95, 487]. Puzzles [410].

qua [167]. Quanta [435]. Quantification [65]. Quantities [406]. Quantization [628]. Quantum [667, 357, 190, 370, 177, 256, 308, 398, 806, 278, 801, 168, 259, 210, 200, 436, 629, 8, 115, 437, 257, 68, 549, 365, 67, 620, 667, 308, 114]. Question [701]. Questions [715]. Quine [295].

Rabbit [746]. Race [151, 337, 338, 364, 216, 336]. Racism [548]. Radical [128, 163, 336]. Randomness [716, 110]. Ranking [139]. Ratchet [418]. Rate [793, 31]. Rate-Distortion [793]. Rates [205]. Rational [627, 377, 227]. Rationality [563]. Ravens [266]. Raynaud [756]. Razor [192, 786, 477]. Razors [488]. Reaction [503]. Reading [394, 808]. Real [588, 75]. Realism [51, 498, 703, 4, 769, 337, 154, 83, 289, 61, 704, 73, 792, 409, 597, 186]. Realism/Antirealism [73]. Reality [330, 486, 115]. Realizability [785]. Realization [24]. Really [663, 185, 19, 662, 545]. Reasonable [97]. Reasoning [577, 399, 594, 721]. Reasons [715, 590]. Reassessing [108, 213]. Rebut [651]. Received [347]. Receiver [323]. Recognition [510]. Recombination [235]. Recommendations [444]. Reconceived [5]. Reconceiving [85]. Reconsidering [114]. Reconstructing [486]. Reconstruction [143]. Recover [434]. Recovering [635]. Recovery [635]. Reduced [800]. Reduction [770, 744, 94, 718]. Reductionism [654, 711, 392, 651, 238]. Redundancy [791]. Redux [197]. Reexamination [306]. Referees [18, 123, 203, 284, 373, 465, 554, 642, 731]. Reference [576, 30, 187, 405, 246, 453, 58, 311]. Reflect [723]. Reflection [180, 199, 133]. Reflections [541]. Regarding [482]. Region [172]. Region- [172]. Regress [221]. Regularities [92, 213]. Regularity [616, 188]. Regulatory [324, 631]. Rehabilitating [789]. Reinforcement [417]. Relation [791, 524]. Relational [404]. Relations [435, 155]. Relationship [635]. Relativistic [278, 68, 67]. Relativity [521, 433, 614, 138, 464, 252]. Relevance [214, 155, 675, 227]. Reliability [155]. Reliable [810, 367]. Remarks [280]. Remedying [779]. Renormalization [636, 703, 437, 345, 704, 346]. Repackaging [106]. Replace [19]. Replicability [600]. Replication [766, 736]. Reply [182, 366, 754, 137, 556, 209, 752]. Report [248]. Reported [428]. Representation [595, 126, 596, 536, 792, 55, 315, 293]. Representational [402, 727]. Representations [563, 793, 78]. Representing [674, 269, 458]. Reproduction [788, 313, 503, 506]. Reproductive [532]. Requirement [570]. Rescuing [133]. Research [659, 147, 679, 611, 672, 352, 501, 163, 164, 319, 681, 274, 173, 630]. Resemblance [330]. Resilience [53]. Resituating [335]. Resolves [314]. Resonance [422]. Resource [472]. Response [601, 87]. Responsible [11]. Responsiveness [603]. Rest [332]. Result [261]. Results [160, 721]. Rethinking [176, 246, 438]. Reuse [324, 671, 709, 791]. Revealing [162]. Reversal [549]. Reversals [745]. Review [540, 667, 656, 488, 357, 273, 359, 383, 718, 371, 184, 741, 370, 122, 308, 577, 382, 114, 755, 130, 183, 307, 296, 384, 294, 566, 104, 666, 309, 210, 717, 578, 295, 139, 486, 271, 367, 297, 11, 164, 442, 154, 369, 115, 138, 640, 358, 9, 103, 193, 140, 396, 485, 283, 553, 395, 281, 368, 455, 567, 282, 394, 456, 10, 655, 641, 729, 293, 730, 272, 474, 541, 487, 756, 456]. Revising [559]. Revisited [7, 379]. Revisiting [733]. Revolution [297]. Revolutions [26, 541]. Reward [600]. Rice [366]. Richard [640]. Richards [541]. Richer [563]. Risk [739, 646, 713, 231, 631, 413, 79, 333, 166, 3, 230, 479]. Road [26]. Robert [541]. Robust [22, 352, 125]. Robustness [603, 609, 159, 63, 166, 557]. Role [142, 334, 147, 234, 683, 55, 110, 492]. Roles [91, 763]. Rosalind [108]. Route [437]. Ruetsche [210]. Rule [320, 57, 509]. Rule-Following [320]. Rules [45, 46, 44, 267]. Ruse [183]. Ryoichi [9].

s [375, 597]. Sabina [656]. Sahotra [283]. Sameness [613, 119]. San [54]. Sandra [103]. Sarkar [283, 356]. Satisfaction [342]. Saunders [115, 451]. Saving [50]. Say [799]. Scale [690, 795]. Schaffner [641]. Scholes [485, 702]. Schupbach [354]. Science [149, 18, 123, 203, 643, 718, 231, 665, 147, 398, 777, 151, 91, 229, 262, 384, 775, 104, 348, 573, 578, 680, 443, 673, 162, 75, 165, 103, 685, 412, 62, 740, 407, 360, 600, 647, 444, 601, 814, 682, 530, 10, 541, 487, 421, 808, 578, 272, 756, 741, 281, 730, 284, 373, 465, 554, 642, 731, 394]. Sciences [804, 512, 611, 131, 736, 670, 740, 720, 242, 491]. Scientific [29, 161, 595, 674, 565, 391, 713, 51, 150, 296, 7, 539, 703, 108, 116, 309, 13, 779, 427, 93, 614, 404, 499, 193, 83, 596, 289, 496, 61, 497, 26, 360, 750, 409, 293, 743, 541, 122, 717, 756]. Scientific/Intellectual [779]. Scientist [11, 671, 167]. Scientists [307, 232]. Scoring [724, 267]. Search [226, 378, 325, 804, 369]. Second [126]. Seeing [810, 367]. Seeks [107]. Segments [725]. Segregation [107]. Selected [562]. Selection [547, 604, 102, 182, 350, 598, 726, 508, 658, 35, 263, 33, 87, 734, 140]. Selective [289, 597]. Semantic [29, 191, 202, 727, 280]. Semantically [72]. Semantics [748]. Semmelweis [251]. Sender [323]. Sender-Receiver [323]. Sense [104, 194, 247]. Sensitivity [197, 411]. Separability [613]. Separating [696]. Sequential [250]. Serious [158]. Setsuko [9]. Seventeenth [238]. Seventeenth-Century [238]. Severe [331]. Sex [680]. Sexuality [307]. Shape [329]. Share [573]. Shared [270]. Sharing [161]. Shattering [251]. Shifting [807, 395]. Should [799, 501, 152, 813, 178, 641]. Signal [48]. Signaling [390, 603, 325, 470, 290, 326]. Significance [222, 452]. Silva [297]. Similarity [760, 158, 485]. Simon [755, 115]. Simple [208]. Simplicity [676]. Simplification [676]. Simulation [485]. Simulations [206, 299]. since [26]. Single [22]. Singularities [339]. Singularity [340]. Size [329]. Skepticism [584, 285]. Sleeping [468]. Small [333, 455]. Smart [563]. Smith [140, 396]. Sober [488, 182]. Social [507, 512, 7, 444, 630]. Societal [665]. Society [272]. Socio [756]. Socio-Historical [756]. Solomonoff [477]. Solution [2, 393, 266, 336, 376, 221]. Solutions [550, 748]. Solvay [114]. Solve [32]. Some [532, 688, 501, 746, 98]. Sometimes [355]. Space [278, 97, 339, 463, 615, 253, 282, 451, 273]. Space-Time [97, 339]. Space-Times [253]. Spaces [96, 688]. Spacetime [81, 705]. Spatially [513]. Spatially-Structurally [513]. Special [595, 380]. Specially [434]. Species [37, 312, 781, 782]. Specificity [385, 761, 564]. Specious [6]. Speech [775, 725, 693]. Spin [655]. Spin-Statistics [655]. Spite [510]. Spohn [139, 752, 751]. Spontaneous [168, 483]. Spots [793]. Sprenger [354]. Springs [386]. Stability [738]. Stabilization [491]. Stable [599, 593, 632]. Stag [449]. Stand [315]. Stand-Ins [315]. Standard [311]. Standards [626, 395, 807]. Stanford [408]. Statement [2]. Statistical [136, 190, 168, 604, 185, 753, 582, 205, 209, 586]. Statistics [655, 145]. Steel [384]. Stellar [73]. Stem [241]. Stephen [122]. Steven [383]. Still [728, 445]. Stochastic [738]. Stotz [294]. Strategies [309, 687, 335, 319, 681, 242, 773, 315]. Strategy [234, 449]. Strength [449]. String [433, 181]. Structural [439]. Structurally [513]. Structure [661, 760, 108, 234, 13, 525, 26, 720, 100, 464, 815, 541, 368]. Structured [653]. Structures [219]. Studies [779, 148, 335]. Study [147, 526, 772, 656]. Studying [307]. Stuff [273]. Style [393, 571, 54]. Subjective [248]. Substance [59]. Substantial [737]. Substrate [270]. Success [532, 499]. Successful [226, 289]. Sufficient [460]. Sugar [447]. Superselection [190]. Supervenience [770]. Supply [638]. Support [286]. Supporting [28]. Supports [180]. Sure [12]. Survival [788]. Survives [86]. Sweet [793]. Symmetries [661, 737, 799, 341, 524, 620]. Symmetry [81, 168, 483, 77, 452, 635]. Syntactic [202]. Synthesizing [665]. Synthetic [235]. System [802, 600, 749]. Systematicity [281]. Systematics [144]. Systems [177, 483, 323, 786, 687, 343, 206, 746, 345, 319, 686, 589, 78]. Szilard [89].

T [9]. Takes [43]. Taking [651]. Tale [171, 681, 784]. Tales [499]. Talk [417, 500]. Tanaka [9]. Tapestry [730]. Targets [289]. Tarski [295]. Task [248]. Taxa [40]. TBD [163]. Teaching [348]. Technique [22]. Teleological [735]. Teleosemantics [318, 666]. Tell [661, 66]. Temperature [100]. Temporal [257, 705]. Tension [77]. Terms [343, 152]. Tested [446]. Testing [344, 609, 413, 247, 31, 223]. Tests [331]. Their [632, 101, 593]. Themes [146]. Theorem [536, 584, 82]. Theorems [126]. Theoretic [517]. Theoretical [191, 381, 596]. Theories [226, 613, 787, 191, 116, 13, 446, 287, 614, 289, 27, 453, 173, 210]. Theory [302, 739, 357, 190, 544, 450, 275, 308, 114, 806, 278, 801, 775, 168, 562, 131, 259, 550, 402, 258, 139, 433, 629, 30, 115, 437, 772, 206, 187, 270, 201, 726, 283, 79, 509, 747, 426, 412, 62, 662, 639, 68, 374, 497, 549, 247, 556, 434, 444, 589, 263, 475, 78, 768, 42, 453, 67, 58, 414, 370, 138]. Theory-Ladenness [412]. Therapies [672]. There [595, 787, 380, 350, 445, 735, 46, 346, 497, 87]. Thermodynamics [56, 261, 58, 622]. Thesis [179, 428, 90]. Thick [72]. Thing [12]. Things [810, 367, 265, 455]. Thinking [682, 144, 311]. Thomas [717]. Thought [683]. Threaten [529]. Three [499, 549, 585]. Thresholds [723]. Thumper [746]. Tickets [776]. Tim [540, 282]. Time [273, 278, 482, 670, 652, 97, 339, 463, 523, 615, 255, 549, 279, 301, 359, 282, 755]. Times [253, 451]. Timescales [670]. Timothy [10]. Timpson [308]. Today [50]. Tolerance [20]. Top [546]. Top-Down [546]. Topics [138]. Topological [688]. Topologists [566]. Topology [253]. Tossing [728]. Total [570, 286]. Toxicology [631]. Tracing [70, 71]. Tracking [689]. Trade [234]. Trade-off [234]. Tragedy [161]. trans [756]. Transaction [756]. Transfer [390]. Transitions [744, 260]. Transitive [460]. Translation [770]. Trautman [635]. Treatment [761]. Tree [144]. Trees [143]. Trials [446, 250]. Triangulation [461]. Trouble [453]. Truism [520].

Trumps [716]. Truth [192, 387]. Truth-Finding [192]. Truths [103]. Tumors [333]. Tuning [570, 530]. Turn [405]. Turning [193]. Twentieth [395, 807]. Twice [391]. Two [507, 391, 126, 154, 764, 784, 280]. Tychomancy [815, 368]. Types [606]. Typicality [169, 617].

UKCP09 [236]. Unacceptable [4]. Uncertainties [65]. Uncertainty [241]. Unconceived [408]. Underdetermination [231, 776, 220]. Undermine [90]. Understand [638]. Understanding [538, 673, 106, 215, 700, 495, 452, 522, 317, 643]. Unfolding [535]. Unification [664, 436, 537, 556]. Unified [292]. Uniformitarianism [409]. Unimaginable [746]. Unit [574]. Units [325, 506]. Unity [500, 35, 291, 491]. Universal [46, 346, 44]. Universality [636, 800, 252]. Universals [271]. Universe [462, 797]. University [667, 640, 641]. Unreal [619]. Unsimple [103]. Updating [300, 129, 224]. Usable [151]. Use [142, 146, 670, 240]. Used [802]. Usefulness [311]. User [488, 358]. Using [546, 405, 601, 479]. Utility [569, 708, 771].

Vacuum [614]. Valentini [114]. Valid [527]. Validation [527, 677]. Validity [445, 796]. Valuable [388]. Value [228, 714, 789, 802, 167, 274, 479, 448]. Value-Free [479]. Values [692, 231, 665, 776, 228, 91, 229, 262, 539, 723, 360, 601, 3, 230, 682, 730, 630, 646]. Variable [95]. Variation [512]. Variations [146]. Varieties [118]. Variety [179]. Variety-of-Evidence [179]. Vectors [326]. Ventriloquism [680]. Veridicality [454]. versus [737, 168, 286, 172, 753, 599, 614, 28, 600, 621, 620]. via [699]. Viability [529]. Vices [600]. View [41, 191, 202, 440, 347, 399, 86, 145, 280]. Views [214]. Vignemont [270]. Vindicated [695]. Violations [218]. Virtue [75]. Virtues [600]. Vision [498]. Visual [576]. Visualization [108]. Vivo [315]. vol [632]. Voles [485]. Volterra [575]. Vulnerable [31].

Wallace [115]. Walsh [25]. War [43]. Warming [66]. Was [74]. Waters [122]. Way [728]. Weak [435]. Weakly [449]. Weatherall [643]. Weisberg [717, 234, 485]. Well [149]. Well-Being [149]. Were [532]. Westermarck [733]. Where [265, 586]. Whewell [43]. Which [519, 425, 564]. Who [178]. Whose [64]. Wilhelm [56]. Will [174]. William [790, 193, 152]. Williamson [728]. Wilson [717]. Wishful [682]. within [323]. without [689, 305, 313, 478, 34, 45, 318, 156, 144, 58]. Wolfgang [139, 752, 751]. Woodward [214, 444, 213, 752]. Words [72]. Work [174, 728]. World [47, 155, 244, 104]. Worlds [387, 115]. Worries [170]. Worrying [500]. Would [683]. Wright [142]. Wrong [39, 33].

xiii [641]. xxii [756].

York [641].

Zealotry [646]. Zeno [306]. Zero [258]. ZFEL [41].

References

Davey:2009:WGC

[1] Kevin Davey. What is Gibbs’s canonical distribution? Philosophy of Science, 76(5):970–983, December 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/ 10.1086/605793. See reply [209].

Crupi:2010:ICS

[2] Vincenzo Crupi and Katya Tentori. Irrelevant conjunction: Statement and solution of a new paradox. Philosophy of Science, 77(1):1–13, January 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/650205.

Steel:2010:EVA

[3] Daniel Steel. Epistemic values and the argument from inductive risk. Philosophy of Science, 77(1):14–34, January 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/650206.

Frost-Arnold:2010:NMA

[4] Greg Frost-Arnold. The no–miracles argument for realism: Inference to an unacceptable explanation. Philosophy of Science, 77(1):35–58, January 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/650207.

Potochnik:2010:LER

[5] Angela Potochnik. Levels of explanation reconceived. Philosophy of Science, 77(1):59–72, January 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/650208.

Barker:2010:SI

[6] Matthew J. Barker. Specious intrinsicalism. Philosophy of Science, 77 (1):73–91, January 2010. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
650209.

Fagan:2010:SCR

[7] Melinda B. Fagan. Social construction revisited: Epistemology and scientific practice. Philosophy of Science, 77(1):92–116, January 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/650210.

Ladyman:2010:PII

[8] James Ladyman and Tomasz Bigaj. The principle of the identity of indiscernibles and quantum mechanics. Philosophy of Science, 77(1):117–136, January 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/650211.

Michael:2010:BRJ

[9] John Michael and Friedrich Stadler. Book review: John T. Blackmore, Ryoichi Itagaki and Setsuko Tanaka (eds.): Ernst Mach’s Philosophy Pro and Con. Philosophy of Science, 77(1):137–140, January 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/650212.

Swan:2010:BRT

[10] Liz Stillwaggon Swan. Book review: Timothy McGrew, Marc Alspector-Kelly and Fritz Allhoff (eds.): Philosophy of Science: An Historical Anthology. Philosophy of Science, 77(1):141–143, January 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/650213.

Kourany:2010:BRJ

[11] Janet Kourany. Book review: John Forge: The Responsible Scientist: A Philosophical Inquiry. Philosophy of Science, 77(1):143–146, January 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/650214.

Walsh:2010:STF

[12] Denis M. Walsh. Not a sure thing: Fitness, probability, and causation. Philosophy of Science, 77(2):147–171, April 2010. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/651320.

Henderson:2010:SDS

[13] Leah Henderson, Noah D. Goodman, Joshua B. Tenenbaum, and James F. Woodward. The structure and dynamics of scientific theories: A hierarchical Bayesian perspective. Philosophy of Science, 77(2):172–200, April 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/651319.

Leitgeb:2010:OJBa

[14] Hannes Leitgeb and Richard Pettigrew. An objective justification of Bayesianism I: Measuring inaccuracy. Philosophy of Science, 77(2):201– 235, April 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/651317.

Leitgeb:2010:OJBb

[15] Hannes Leitgeb and Richard Pettigrew. An objective justification of Bayesianism II: The consequences of minimizing inaccuracy. Philosophy of Science, 77(2):236–272, April 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/651318.

Weslake:2010:ED

[16] Brad Weslake. Explanatory depth. Philosophy of Science, 77(2):273– 294, April 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/651316.

Kulvicki:2010:KIM

[17] John Kulvicki. Knowing with images: Medium and message. Philosophy of Science, 77(2):295–313, April 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/651321.

Anonymous:2010:RPS

[18] Anonymous. Referees for philosophy of science. Philosophy of Science, 77(2):314, April 2010. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
651730.

Leuridan:2010:CMR

[19] Bert Leuridan. Can mechanisms really replace laws of nature? Philosophy of Science, 77(3):317–340, July 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/652959.

Hudson:2010:CPT

[20] Robert Hudson. Carnap, the principle of tolerance, and empiricism. Philosophy of Science, 77(3):341–358, July 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/652960.

Spanos:2010:DAC

[21] Aris Spanos. The discovery of argon: A case for learning from data? Philosophy of Science, 77(3):359–380, July 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/652961.

Gandenberger:2010:PRB

[22] Gregory Stephen Gandenberger. Producing a robust body of data with a single technique. Philosophy of Science, 77(3):381–399, July 2010. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/652962.

Shapiro:2010:JBB

[23] Lawrence Shapiro. James bond and the barking dog: Evolution and extended cognition. Philosophy of Science, 77(3):400–418, July 2010. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/652963.

Figdor:2010:NMR

[24] Carrie Figdor. Neuroscience and the multiple realization of cognitive functions. Philosophy of Science, 77(3):419–456, July 2010. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/652964.

Northcott:2010:WCE

[25] Robert Northcott. Walsh on causes and evolution. Philosophy of Science, 77(3):457–467, July 2010. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
652965.

Reydon:2010:DKE

[26] Thomas A. C. Reydon and Paul Hoyningen-Huene. Discussion: Kuhn’s evolutionary analogy in the structure of scientific revolutions and “The Road since Structure”. Philosophy of Science, 77(3):468–476, July 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/652966.

Shagrir:2010:MCL

[27] Oron Shagrir. Marr on computational-level theories. Philosophy of Science, 77(4):477–500, October 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/656005.

Norton:2010:CCS

[28] John D. Norton. Cosmic confusions: Not supporting versus supporting not. Philosophy of Science, 77(4):501–523, October 2010. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/661504.

Andreas:2010:SHS

[29] Holger Andreas. Semantic holism in scientific language. Philosophy of Science, 77(4):524–543, October 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/656007.

Kuukkanen:2010:KEC

[30] Jouni-Matti Kuukkanen. Kuhn on essentialism and the causal theory of reference. Philosophy of Science, 77(4):544–564, October 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/656008.

Spanos:2010:FTV

[31] Aris Spanos. Is frequentist testing vulnerable to the base-rate fallacy? Philosophy of Science, 77(4):565–583, October 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/656009.

Arnold:2010:CBA

[32] Eckhart Arnold. Can the best-alternative justification solve Hume’s problem? On the limits of a promising approach. Philosophy of Science, 77 (4):584–593, October 2010. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
656010.

Sober:2010:NSC

[33] Elliott Sober. Natural selection, causality, and laws: What Fodor and Piatelli-Palmarini got wrong. Philosophy of Science, 77(4):594–607, October 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/10.1086/656020.

Gildenhuys:2010:CEC

[34] Peter Gildenhuys. Causal equations without Ceteris Paribus clauses. Philosophy of Science, 77(4):608–632, October 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/656021.

Roszkowski:2010:NSU

[35] Adela Helena Roszkowski. Natural selection and the unity of functional analyses. Philosophy of Science, 77(4):633–645, October 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/656022.

Richardson:2010:P

[36] Alan Richardson. Preface. Philosophy of Science, 77(5):647, December 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/657990.

Devitt:2010:SPI

[37] Michael Devitt. Species have (partly) intrinsic essences. Philosophy of Science, 77(5):648–661, December 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/656820.

Rieppel:2010:NEB

[38] Olivier Rieppel. New essentialism in biology. Philosophy of Science, 77 (5):662–673, December 2010. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
656539.

Ereshefsky:2010:WWN

[39] Marc Ereshefsky. What’s wrong with the new biological essentialism. Philosophy of Science, 77(5):674–685, December 2010. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/656545.

Boyd:2010:HHT

[40] Richard Boyd. Homeostasis, higher taxa, and monophyly. Philosophy of Science, 77(5):686–701, December 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/656551.

Brandon:2010:NNN

[41] Robert N. Brandon. A non-Newtonian Newtonian model of evolution: The ZFEL view. Philosophy of Science, 77(5):702–715, December 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/656901.

Stephens:2010:FCE

[42] Christopher Stephens. Forces and causes in evolutionary theory. Philosophy of Science, 77(5):716–727, December 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/656821.

Achinstein:2010:WIW

[43] Peter Achinstein. The War on induction: Whewell takes on Newton and mill (norton takes on everyone). Philosophy of Science, 77(5):728–739, December 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/656540.

Worrall:2010:URA

[44] John Worrall. For universal rules, against induction. Philosophy of Science, 77(5):740–753, December 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/656823.

Kelly:2010:HNI

[45] Thomas Kelly. Hume, norton, and induction without rules. Philosophy of Science, 77(5):754–764, December 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/656548.

Norton:2010:TNU

[46] John D. Norton. There are no universal rules for induction. Philosophy of Science, 77(5):765–777, December 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/656542.

Bogen:2010:NW

[47] Jim Bogen. Noise in the world. Philosophy of Science, 77(5):778–791, December 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/656006.

Woodward:2010:DPS

[48] James Woodward. Data, phenomena, signal, and noise. Philosophy of Science, 77(5):792–803, December 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/656554.

McAllister:2010:OPE

[49] James W. McAllister. The ontology of patterns in empirical data. Philosophy of Science, 77(5):804–814, December 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/656555.

Teller:2010:SPT

[50] Paul Teller. “Saving the phenomena” today. Philosophy of Science, 77 (5):815–826, December 2010. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
656537.

Brading:2010:APS

[51] Katherine Brading. Autonomous patterns and scientific realism. Philosophy of Science, 77(5):827–839, December 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/656816.

Craver:2010:PM

[52] Carl F. Craver. Prosthetic models. Philosophy of Science, 77(5):840–851, December 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/656822.

Piccinini:2010:RC

[53] Gualtiero Piccinini. The resilience of computationalism. Philosophy of Science, 77(5):852–861, December 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/656549.

Shagrir:2010:CSD

[54] Oron Shagrir. Computation, San Diego style. Philosophy of Science, 77 (5):862–874, December 2010. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
656553.

Sullivan:2010:RRC

[55] Jacqueline Anne Sullivan. A role for representation in cognitive neurobiology. Philosophy of Science, 77(5):875–887, December 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/656818.

Deltete:2010:TWO

[56] Robert J. Deltete. Thermodynamics in wilhelm Ostwald’s physical chemistry. Philosophy of Science, 77(5):888–899, December 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/656824.

Needham:2010:MIP

[57] Paul Needham. A mereological interpretation of the phase rule. Philosophy of Science, 77(5):900–910, December 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/656546.

Vemulapalli:2010:TCH

[58] G. K. Vemulapalli. Thermodynamics and chemistry: How does a theory formulated without reference to matter explain the properties of matter? Philosophy of Science, 77(5):911–920, December 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/657481.

Hendry:2010:ECS

[59] Robin Findlay Hendry. Entropy and chemical substance. Philosophy of Science, 77(5):921–932, December 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/656550.

Wilson:2010:MLE

[60] Mark Wilson. Mixed-level explanation. Philosophy of Science, 77(5): 933–946, December 2010. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
657426.

Psillos:2010:SRB

[61] Stathis Psillos. Scientific realism: Between Platonism and nominalism. Philosophy of Science, 77(5):947–958, December 2010. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/656825.

Pincock:2010:MSC

[62] Christopher Pincock. Mathematics, science, and confirmation theory. Philosophy of Science, 77(5):959–970, December 2010. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/656552.

Lloyd:2010:CRC

[63] Elisabeth A. Lloyd. Confirmation and robustness of climate models. Philosophy of Science, 77(5):971–984, December 2010. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/657427.

Parker:2010:WPP

[64] Wendy S. Parker. Whose probabilities? Predicting climate change with ensembles of models. Philosophy of Science, 77(5):985–997, December 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/656815.

Mearns:2010:QUF

[65] Linda O. Mearns. Quantification of uncertainties of future climate change: Challenges and applications. Philosophy of Science, 77(5):998–1011, December 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/656817.

Oreskes:2010:AGW

[66] Naomi Oreskes, David A. Stainforth, and Leonard A. Smith. Adaptation to global warming: Do climate models tell us what we need to know? Philosophy of Science, 77(5):1012–1028, December 2010. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/657428.

Valente:2010:CED

[67] Giovanni Valente. Can entanglement be destroyed by any local operation in relativistic quantum field theory? Philosophy of Science, 77(5): 1029–1041, December 2010. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
656547.

Redei:2010:EDN

[68] Miklós Rédei. Einstein’s dissatisfaction with nonrelativistic quantum mechanics and relativistic quantum field theory. Philosophy of Science, 77(5): 1042–1057, December 2010. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
656819.

Steel:2010:NAA

[69] Daniel Steel. A new approach to argument by analogy: Extrapolation and chain graphs. Philosophy of Science, 77(5):1058–1069, December 2010.

CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/656543.

Guala:2010:EAC

[70] Francesco Guala. Extrapolation, analogy, and comparative process tracing. Philosophy of Science, 77(5):1070–1082, December 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/656541.

Parker:2010:CPT

[71] Wendy S. Parker. Comparative process tracing and climate change fingerprints. Philosophy of Science, 77(5):1083–1095, December 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/656814.

Werning:2010:CFE

[72] Markus Werning. Complex first? On the evolutionary and developmental priority of semantically thick words. Philosophy of Science, 77(5): 1096–1108, December 2010. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
656826.

Ruphy:2010:SKN

[73] Stéphanie Ruphy. Are stellar kinds natural kinds? A challenging newcomer in the monism/pluralism and realism/antirealism debates. Philosophy of Science, 77(5):1109–1120, December 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/656544.

Sklar:2010:ILN

[74] Lawrence Sklar. I’d Love to be a naturalist — if only I knew what naturalism was. Philosophy of Science, 77(5):1121–1137, December 2010. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/656827.

Korkut:2011:RVF

[75] Buket Korkut. The real virtue of Friedman’s neo-Kantian philosophy of science. Philosophy of Science, 78(1):1–15, January 2011. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/658092.

Davey:2011:ICP

[76] Kevin Davey. Idealizations and contextualism in physics. Philosophy of Science, 78(1):16–38, January 2011. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/658093.

Jantzen:2011:AST

[77] Benjamin Jantzen. An awkward symmetry: The tension between particle ontologies and permutation invariance. Philosophy of Science, 78(1):39– 59, January 2011. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/658094.

Shea:2011:DST

[78] Nicholas Shea. Developmental systems theory formulated as a claim about inherited representations. Philosophy of Science, 78(1):60–82, January 2011. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/658110.

Okasha:2011:OCF

[79] Samir Okasha. Optimal choice in the face of risk: Decision theory meets evolution. Philosophy of Science, 78(1):83–104, January 2011. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/658115.

Schupbach:2011:LEP

[80] Jonah N. Schupbach and Jan Sprenger. The logic of explanatory power. Philosophy of Science, 78(1):105–127, January 2011. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.

jstor.org/stable/10.1086/658111.

Baker:2011:BSS

[81] David John Baker. Broken symmetry and spacetime. Philosophy of Science, 78(1):128–148, January 2011. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/658112.

Ainsworth:2011:WCD

[82] Peter M. Ainsworth. What chains does Liouville’s Theorem put on Maxwell’s Demon? Philosophy of Science, 78(1):149–164, January 2011. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/658113.

Neuber:2011:FSR

[83] Matthias Neuber. Feigl’s ‘scientific realism’. Philosophy of Science, 78 (1):165–183, January 2011. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
658114.

Anonymous:2011:E

[84] Anonymous. Erratum. Philosophy of Science, 78(1):184, January 2011. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/658274.

Forber:2011:REI

[85] Patrick Forber. Reconceiving eliminative inference. Philosophy of Science, 78(2):185–208, April 2011. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
659232.

Otsuka:2011:WCV

[86] Jun Otsuka, Trin Turner, Colin Allen, and Elisabeth A. Lloyd. Why the causal view of fitness survives. Philosophy of Science, 78(2):209–224, April 2011. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/659219.

Weinberger:2011:TED

[87] Naftali Weinberger. Is there an empirical disagreement between genic and genotypic selection models? A response to Brandon and Nijhout. Philosophy of Science, 78(2):225–237, April 2011. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/659220.

Zednik:2011:NDE

[88] Carlos Zednik. The nature of dynamical explanation. Philosophy of Science, 78(2):238–263, April 2011. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/659221.

Hemmo:2011:SPM

[89] Meir Hemmo and Orly Shenker. Szilard’s Perpetuum Mobile. Philosophy of Science, 78(2):264–283, April 2011. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/659231.

Moffatt:2011:CCA

[90] Barton Moffatt. Conflations in the causal account of information undermine the parity thesis. Philosophy of Science, 78(2):284–302, April 2011. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/659230.

Elliott:2011:DIR

[91] Kevin C. Elliott. Direct and indirect roles for values in science. Philosophy of Science, 78(2):303–324, April 2011. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/659222.

Andersen:2011:MLR

[92] Holly K. Andersen. Mechanisms, laws, and regularities. Philosophy of Science, 78(2):325–331, April 2011. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/659229.

Lange:2011:CLS

[93] Marc Lange. Conservation laws in scientific explanations: Constraints or coincidences? Philosophy of Science, 78(3):333–352, July 2011. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/660299. See critique [124].

vanRiel:2011:NRB

[94] Raphael van Riel. Nagelian reduction beyond the Nagel model. Philosophy of Science, 78(3):353–375, July 2011. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/660300.

Schlosshauer:2011:FCC

[95] Maximilian Schlosshauer and Gregory Wheeler. Focused correlation, confirmation, and the jigsaw puzzle of variable evidence. Philosophy of Science, 78(3):376–392, July 2011. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/659218.

Gyenis:2011:CCC

[96] Zalán Gyenis and Miklós Rédei. Characterizing common cause closed probability spaces. Philosophy of Science, 78(3):393–409, July 2011. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/660302.

Manchak:2011:WPR

[97] John Byron Manchak. What is a physically reasonable space-time? Philosophy of Science, 78(3):410–420, July 2011. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/660301.

Weatherall:2011:SEP

[98] James Owen Weatherall. On (some) explanations in physics. Philosophy of Science, 78(3):421–447, July 2011. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/660737.

Glymour:2011:MEI

[99] Bruce Glymour. Modeling environments: Interactive causation and adaptations to environmental conditions. Philosophy of Science, 78(3):448–471, July 2011. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/10.1086/660303.

Skow:2011:DTM

[100] Bradford Skow. Does temperature have a metric structure? Philosophy of Science, 78(3):472–489, July 2011. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/660304.

Martinez:2011:EGA

[101] Sergio F. Martínez and Xiang Huang. Epistemic groundings of abstraction and their cognitive dimension. Philosophy of Science, 78(3):490–511, July 2011. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/660305.

Havstad:2011:DPN

[102] Joyce C. Havstad. Discussion: Problems for natural selection as a mechanism. Philosophy of Science, 78(3):512–523, July 2011. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/660734.

Mikkelson:2011:BRS

[103] Gregory M. Mikkelson. Book review: Sandra D. Mitchell, Unsimple Truths: Science, Complexity, and Policy. Philosophy of Science, 78 (3):524–527, July 2011. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
660735.

Garber:2011:BRP

[104] Daniel Garber. Book review: Peter Dear, The Intelligibility of Nature: How Science Makes Sense of the World. Philosophy of Science, 78(3):527– 531, July 2011. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/660736.

Hunt:2012:AHH

[105] J. Christopher Hunt. On ad hoc hypotheses. Philosophy of Science, 79(1):1–14, January 2012. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
663238.

Khalifa:2012:IUR

[106] Kareem Khalifa. Inaugurating understanding or repackaging explanation? Philosophy of Science, 79(1):15–37, January 2012. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/663235.

Muldoon:2012:SNO

[107] Ryan Muldoon, Tony Smith, and Michael Weisberg. Segregation that no one seeks. Philosophy of Science, 79(1):38–62, January 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/663236.

Gibbons:2012:RDR

[108] Michelle G. Gibbons. Reassessing discovery: Rosalind Franklin, scientific visualization, and the structure of DNA. Philosophy of Science, 79(1):63– 80, January 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/663241.

Wagner:2012:EDF

[109] Elliott O. Wagner. Evolving to divide the fruits of cooperation. Philosophy of Science, 79(1):81–94, January 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/663243.

Wagner:2012:RRD

[110] Andreas Wagner. The role of randomness in Darwinian evolution. Philosophy of Science, 79(1):95–119, January 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/663239.

Potochnik:2012:LHO

[111] Angela Potochnik and Brian McGill. The limitations of hierarchical organization. Philosophy of Science, 79(1):120–140, January 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/663237.

Morrison:2012:EPM

[112] Margaret Morrison. Emergent physics and micro-ontology. Philosophy of Science, 79(1):141–166, January 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/663240.

Fitelson:2012:ALD

[113] Branden Fitelson. Accuracy, language dependence, and Joyce’s argument for probabilism. Philosophy of Science, 79(1):167–174, January 2012. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/663242. See reply [137].

Dickson:2012:BRG

[114] Michael Dickson. Book review: Guido Bacciagaluppi and Antony Valentini, Quantum Theory at the Crossroads: Reconsidering the 1927 Solvay Conference. Philosophy of Science, 79(1):175–177, January 2012. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/663245.

Lewis:2012:BRS

[115] Peter Lewis. Book review: Simon Saunders, Jonathan Barrett, Adrian Kent, and David Wallace, Many Worlds? Everett, Quantum Theory, and Reality. Philosophy of Science, 79(1):177–181, January 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/663244.

Halvorson:2012:WST

[116] Hans Halvorson. What scientific theories could not be. Philosophy of Science, 79(2):183–206, April 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/664745.

Norton:2012:AIW

[117] John D. Norton. Approximation and idealization: Why the difference matters. Philosophy of Science, 79(2):207–232, April 2012. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/664746.

Nathan:2012:VME

[118] Marco J. Nathan. The varieties of molecular explanation. Philosophy of Science, 79(2):233–254, April 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/664791.

Ramsey:2012:SB

[119] Grant Ramsey and Anne Siebels Peterson. Sameness in biology. Philosophy of Science, 79(2):255–275, April 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/664744.

Baetu:2012:MCE

[120] Tudor M. Baetu. Mechanistic constraints on evolutionary outcomes. Philosophy of Science, 79(2):276–294, April 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/664743.

deVignemont:2012:WIL

[121] Frédérique de Vignemont and Pierre Jacob. What is it like to feel another’s pain? Philosophy of Science, 79(2):295–316, April 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/664742. See discussion [270].

Cat:2012:BRS

[122] Jordi Cat. Book review: Stephen H. Kellert, Helen E. Longino, and C. Kenneth Waters, Scientific Pluralism. Philosophy of Science, 79(2):317– 325, April 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/664747.

Anonymous:2012:RPS

[123] Anonymous. Referees for philosophy of science. Philosophy of Science, 79(2):326–331, April 2012. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
665592.

Demarest:2012:DCG

[124] Heather Demarest. Do counterfactuals ground the laws of nature? A critique of Lange. Philosophy of Science, 79(3):333–344, July 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/666066. See [93].

Houkes:2012:RHC

[125] Wybo Houkes and Krist Vaesen. Robust! Handle with care. Philosophy of Science, 79(3):345–364, July 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/666061.

Crupi:2012:SLL

[126] Vincenzo Crupi and Katya Tentori. A second look at the logic of explanatory power (with two novel representation theorems). Philosophy of Science, 79(3):365–385, July 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/666063.

Ahmed:2012:PB

[127] Arif Ahmed. Push the button. Philosophy of Science, 79(3):386–395, July 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/666065.

Earnshaw-Whyte:2012:IRC

[128] Eugene Earnshaw-Whyte. Increasingly radical claims about heredity and fitness. Philosophy of Science, 79(3):396–412, July 2012. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/666060.

Levinstein:2012:LPA

[129] Benjamin Anders Levinstein. Leitgeb and Pettigrew on accuracy and updating. Philosophy of Science, 79(3):413–424, July 2012. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/666064.

Douglas:2012:BRK

[130] Heather Douglas. Book review: Kevin Elliott, Is a Little Pollution Good for You? Philosophy of Science, 79(3):425–428, July 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/666062.

Goodwin:2012:ETP

[131] William Goodwin. Experiments and theory in the preparative sciences. Philosophy of Science, 79(4):429–447, October 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.

jstor.org/stable/10.1086/668003.

Fagan:2012:JAM

[132] Melinda Bonnie Fagan. The joint account of mechanistic explanation. Philosophy of Science, 79(4):448–472, October 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/668006.

Park:2012:RR

[133] Ilho Park. Rescuing reflection. Philosophy of Science, 79(4):473–489, October 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/668001.

Machery:2012:DNC

[134] Edouard Machery. Dissociations in neuropsychology and cognitive neuroscience. Philosophy of Science, 79(4):490–518, October 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/668002.

Hausman:2012:HNF

[135] Daniel M. Hausman. Health, naturalism, and functional efficiency. Philosophy of Science, 79(4):519–541, October 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/668005.

Ainsworth:2012:GPD

[136] Peter M. Ainsworth. The Gibbs Paradox and the definition of entropy in statistical mechanics. Philosophy of Science, 79(4):542–560, October 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/668004.

Pedersen:2012:WLD

[137] Arthur Paul Pedersen and Clark Glymour. What language dependence problem? A reply for Joyce to Fitelson on Joyce. Philosophy of Science, 79(4):561–574, October 2012. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/

667997. See [113].

Manchak:2012:BRD

[138] John Byron Manchak. Book review: David Malament, Topics in the Foundations of General Relativity and Newtonian Gravitation Theory. Philosophy of Science, 79(4):575–583, October 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/668000.

Huber:2012:BRW

[139] Franz Huber. Book review: Wolfgang Spohn, The Laws of Belief: Ranking Theory and Its Philosophical Applications. Philosophy of Science, 79 (4):584–588, October 2012. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/

667998. See comments [751, 752].

OConnor:2012:BRP

[140] Cailin O’Connor. Book review: Peter Godfrey-Smith, Darwinian Populations and Natural Selection. Philosophy of Science, 79(4):589–593, October 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/667999.

Downes:2012:P

[141] Stephen M. Downes. Preface. Philosophy of Science, 79(5):595, December 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/668040.

Abrams:2012:IUW

[142] Marshall Abrams. Implications of use of Wright’s F_{ST} for the role of probability and causation in evolution. Philosophy of Science, 79(5):596–608, December 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/667905.

Haber:2012:MLM

[143] Matthew H. Haber. Multilevel lineages and multidimensional trees: The levels of lineage and phylogeny reconstruction. Philosophy of Science, 79 (5):609–623, December 2012. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
667849.

Velasco:2012:FST

[144] Joel D. Velasco. The future of systematics: Tree thinking without the tree. Philosophy of Science, 79(5):624–636, December 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/667878.

Weir:2012:ESH

[145] Bruce S. Weir. Estimating F -statistics: A historical view. Philosophy of Science, 79(5):637–643, December 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/667904.

Ankeny:2012:DTV

[146] Rachel A. Ankeny. Detecting themes and variations: The use of cases in developmental biology. Philosophy of Science, 79(5):644–654, December 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/667903.

Crasnow:2012:RCS

[147] Sharon Crasnow. The role of case study research in political science: Evidence for causal claims. Philosophy of Science, 79(5):655–666, December 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/667869.

Morgan:2012:CSO

[148] Mary S. Morgan. Case studies: One observation or many? Justification or discovery? Philosophy of Science, 79(5):667–677, December 2012. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/667848.

Alexandrova:2012:WBO

[149] Anna Alexandrova. Well-being as an object of science. Philosophy of Science, 79(5):678–689, December 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/667870.

Chang:2012:APE

[150] Hasok Chang. Acidity: The persistence of the everyday in the scientific. Philosophy of Science, 79(5):690–700, December 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/667900.

Efstathiou:2012:HOR

[151] Sophia Efstathiou. How ordinary race concepts get to be usable in biomedical science: An account of founded race concepts. Philosophy of Science, 79(5):701–713, December 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/667901.

Radick:2012:SHI

[152] Gregory Radick. Should “heredity” and “inheritance” be biological terms? William Bateson’s change of mind as a historical and philosophical problem. Philosophy of Science, 79(5):714–724, December 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/667872.

Bokulich:2012:DEN

[153] Alisa Bokulich. Distinguishing explanatory from nonexplanatory fictions. Philosophy of Science, 79(5):725–737, December 2012. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/667991.

Levy:2012:BRB

[154] Arnon Levy. Book review: Models, Fictions, and Realism: Two Packages. Philosophy of Science, 79(5):738–748, December 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/667992.

Peschard:2012:FMW

[155] Isabelle Peschard. Forging model/world relations:relevance and reliability. Philosophy of Science, 79(5):749–760, December 2012. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/667846.

Thomson-Jones:2012:MM

[156] Martin Thomson-Jones. Modeling without mathematics. Philosophy of Science, 79(5):761–772, December 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/667876.

vanFraassen:2012:MMC

[157] Bas C. van Fraassen. Modeling and measurement: The criterion of empirical grounding. Philosophy of Science, 79(5):773–784, December 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/667847.

Weisberg:2012:GSA

[158] Michael Weisberg. Getting serious about similarity. Philosophy of Science, 79(5):785–794, December 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/667845.

Justus:2012:EBI

[159] James Justus. The elusive basis of inferential robustness. Philosophy of Science, 79(5):795–807, December 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/667902.

Machery:2012:PNR

[160] Edouard Machery. Power and negative results. Philosophy of Science, 79 (5):808–820, December 2012. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
667877.

Biddle:2012:TAI

[161] Justin B. Biddle. Tragedy of the anticommons? Intellectual property and the sharing of scientific information. Philosophy of Science, 79(5):821–832, December 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/667874.

Knuuttila:2012:CCR

[162] Tarja Knuuttila. Contradictions of commercialization: Revealing the norms of science? Philosophy of Science, 79(5):833–844, December 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/667844.

Kukla:2012:ATR

[163] Rebecca Kukla. “author TBD”: Radical collaboration in contemporary biomedical research. Philosophy of Science, 79(5):845–858, December 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/668042.

Lee:2012:KCP

[164] Carole J. Lee. A Kuhnian critique of psychometric research on peer review. Philosophy of Science, 79(5):859–870, December 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/667841.

McAllister:2012:CSC

[165] James W. McAllister. Climate science controversies and the demand for access to empirical data. Philosophy of Science, 79(5):871–880, December 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/667871.

Sprenger:2012:ERA

[166] Jan Sprenger. Environmental risk analysis: Robustness is essential for precaution. Philosophy of Science, 79(5):881–892, December 2012. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/667873.

Steele:2012:SQP

[167] Katie Steele. The scientist qua policy advisor makes value judgments. Philosophy of Science, 79(5):893–904, December 2012. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/667842.

Fraser:2012:SSB

[168] Doreen Fraser. Spontaneous symmetry breaking: Quantum statistical mechanics versus quantum field theory. Philosophy of Science, 79(5):905–916, December 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/668044.

Frigg:2012:DT

[169] Roman Frigg and Charlotte Werndl. Demystifying typicality. Philosophy of Science, 79(5):917–929, December 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/668043.

Stoltzner:2012:CHM

[170] Michael Stöltzner. Constraining the Higgs mechanism: Ontological worries and the prospects for an algebraic cure. Philosophy of Science, 79 (5):930–941, December 2012. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
667875.

Hitchcock:2012:PCD

[171] Christopher Hitchcock. Portable causal dependence: A tale of consilience. Philosophy of Science, 79(5):942–951, December 2012. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/667899.

Klein:2012:COR

[172] Colin Klein. Cognitive ontology and region- versus network-oriented analyses. Philosophy of Science, 79(5):952–960, December 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/667843.

Woodward:2012:CIB

[173] James Woodward. Causation: Interactions between philosophical theories and psychological research. Philosophy of Science, 79(5):961–972, December 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/667850.

Cartwright:2012:WPW

[174] Nancy Cartwright. Will this policy work for you? Predicting effectiveness better: How philosophy helps. Philosophy of Science, 79(5):973–989, December 2012. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/668041.

Rathkopf:2013:LIF

[175] Charles A. Rathkopf. Localization and intrinsic function. Philosophy of Science, 80(1):1–21, January 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/668878.

Saunders:2013:RNP

[176] Simon Saunders. Rethinking Newton’s Principia. Philosophy of Science, 80(1):22–48, January 2013. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
668881.

Caulton:2013:DIQ

[177] Adam Caulton. Discerning “indistinguishable” quantum systems. Philosophy of Science, 80(1):49–72, January 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/668874.

Spanos:2013:WSA

[178] Aris Spanos. Who should be afraid of the Jeffreys–Lindley Paradox? Philosophy of Science, 80(1):73–93, January 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/668875.

Claveau:2013:ICV

[179] François Claveau. The Independence condition in the variety-of-evidence thesis. Philosophy of Science, 80(1):94–118, January 2013. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/668877.

Easwaran:2013:EAS

[180] Kenny Easwaran. Expected accuracy supports conditionalization — and conglomerability and reflection. Philosophy of Science, 80(1):119–142, January 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/668879.

Laraudogoitia:2013:NSP

[181] Jon Pérez Laraudogoitia. The new string paradox. Philosophy of Science, 80(1):143–154, January 2013. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
668876.

Helgeson:2013:WSC

[182] Casey Helgeson. What selection can and cannot explain: A reply to Nanay’s critique of Sober. Philosophy of Science, 80(1):155–159, January 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/668880.

Driscoll:2013:BRM

[183] Catherine Driscoll. Book review: Michael Ruse, The Philosophy of Human Evolution. Philosophy of Science, 80(1):160–164, January 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/668883.

Broadbent:2013:BRJ

[184] Alex Broadbent. Book review: Jeremy Howick, The Philosophy of Evidence-Based Medicine. Philosophy of Science, 80(1):165–168, January 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/668882.

Lange:2013:RSE

[185] Marc Lange. Really statistical explanations and genetic drift. Philosophy of Science, 80(2):169–188, April 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/670323.

Vickers:2013:CCR

[186] Peter Vickers. A confrontation of convergent realism. Philosophy of Science, 80(2):189–211, April 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/670297.

Michael:2013:ACT

[187] John Michael and Miles MacLeod. Applying the causal theory of reference to intentional concepts. Philosophy of Science, 80(2):212–230, April 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/670328.

Pruss:2013:PRC

[188] Alexander R. Pruss. Probability, regularity, and cardinality. Philosophy of Science, 80(2):231–240, April 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/670299.

Levy:2013:AOM

[189] Arnon Levy and William Bechtel. Abstraction and the organization of mechanisms. Philosophy of Science, 80(2):241–261, April 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/670300.

Baker:2013:IST

[190] David John Baker. Identity, superselection theory, and the statistical properties of quantum fields. Philosophy of Science, 80(2):262–285, April 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/670296.

Glymour:2013:TES

[191] Clark Glymour. Theoretical equivalence and the semantic view of theories. Philosophy of Science, 80(2):286–297, April 2013. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/670261.

Fitzpatrick:2013:KOR

[192] Simon Fitzpatrick. Kelly on Ockham’s razor and truth-finding efficiency. Philosophy of Science, 80(2):298–309, April 2013. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/670298.

Miyake:2013:BRW

[193] Teru Miyake. Book review: William Harper, Isaac Newton’s Scientific Method: Turning Data into Evidence about Gravity and Cosmology. Philosophy of Science, 80(2):310–316, April 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/670660.

Garson:2013:FSM

[194] Justin Garson. The functional sense of mechanism. Philosophy of Science, 80(3):317–333, July 2013. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
671173.

Rohwer:2013:HPI

[195] Yasha Rohwer and Collin Rice. Hypothetical pattern idealization and explanatory models. Philosophy of Science, 80(3):334–355, July 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/671399.

Colombo:2013:MFB

[196] Matteo Colombo. Moving forward (and beyond) the modularity debate: A network perspective. Philosophy of Science, 80(3):356–377, July 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/670331.

Brossel:2013:PMS

[197] Peter Brössel. The problem of measure sensitivity redux. Philosophy of Science, 80(3):378–397, July 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/671172.

Rizza:2013:AMB

[198] Davide Rizza. The applicability of mathematics: Beyond mapping accounts. Philosophy of Science, 80(3):398–412, July 2013. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/670295.

Huttegger:2013:DR

[199] Simon M. Huttegger. In defense of reflection. Philosophy of Science, 80(3):413–433, July 2013. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
671427.

Healey:2013:OQO

[200] Richard Healey. Observation and quantum objectivity. Philosophy of Science, 80(3):434–453, July 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/671106.

Ney:2013:DAP

[201] Alyssa Ney and Kathryn Phillips. Does an adequate physical theory demand a primitive ontology? Philosophy of Science, 80(3):454–474, July 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/671076.

Halvorson:2013:SVI

[202] Hans Halvorson. The semantic view, if plausible, is syntactic. Philosophy of Science, 80(3):475–478, July 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/671077.

Anonymous:2013:RPS

[203] Anonymous. Referees for philosophy of science. Philosophy of Science, 80(3):479–482, July 2013. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
672158.

Belot:2013:BO

[204] Gordon Belot. Bayesian orgulity. Philosophy of Science, 80(4):483–503, October 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/673249.

Schervish:2013:EER

[205] Mark J. Schervish, Teddy Seidenfeld, and Joseph B. Kadane. The effect of exchange rates on statistical decisions. Philosophy of Science, 80(4):504– 532, October 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/673267.

MacLeod:2013:BSG

[206] Miles MacLeod and Nancy J. Nersessian. Building simulations from the ground up: Modeling and theory in systems biology. Philosophy of Science, 80(4):533–556, October 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/673209.

Kochiras:2013:MPN

[207] Hylarie Kochiras. The mechanical philosophy and Newton’s mechanical force. Philosophy of Science, 80(4):557–578, October 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/673210.

Roberts:2013:SFC

[208] Bryan W. Roberts. The simple failure of Curie’s Principle. Philosophy of Science, 80(4):579–592, October 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/673212.

Shech:2013:GBR

[209] Elay Shech. On gases in boxes: A reply to Davey on the justification of the probability measure in Boltzmannian statistical mechanics. Philosophy of Science, 80(4):593–605, October 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/673211. See [1].

Healey:2013:BRL

[210] Richard Healey. Book review: Laura Ruetsche, Interpreting Quantum Theories. Philosophy of Science, 80(4):606–609, October 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/673217.

Anonymous:2013:E

[211] Anonymous. Erratum. Philosophy of Science, 80(4):610, October 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/673382.

Woody:2013:P

[212] Andrea I. Woody. Preface. Philosophy of Science, 80(5):611–612, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/674136.

Saatsi:2013:RWA

[213] Juha Saatsi and Mark Pexton. Reassessing Woodward’s account of explanation: Regularities, counterfactuals, and noncausal explanations. Philosophy of Science, 80(5):613–624, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/673899.

Imbert:2013:RIE

[214] Cyrille Imbert. Relevance, not invariance, explanatoriness, not manipulability: Discussion of Woodward’s views on explanatory relevance. Philosophy of Science, 80(5):625–636, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/674002.

Kuorikoski:2013:UNF

[215] Jaakko Kuorikoski and Samuli Pöyhönen. Understanding nonmodular functionality: Lessons from genetic algorithms. Philosophy of Science, 80 (5):637–649, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
673866.

Marcellesi:2013:RC

[216] Alexandre Marcellesi. Is race a cause? Philosophy of Science, 80(5): 650–659, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
673721.

Zwier:2013:ECI

[217] Karen R. Zwier. An epistemology of causal inference from experiment. Philosophy of Science, 80(5):660–671, December 2013. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/673901.

Andersen:2013:WEV

[218] Holly Andersen. When to expect violations of causal faithfulness and why it matters. Philosophy of Science, 80(5):672–683, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/673937.

Eberhardt:2013:EIC

[219] Frederick Eberhardt. Experimental indistinguishability of causal structures. Philosophy of Science, 80(5):684–696, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/673865.

Miyake:2013:UBB

[220] Teru Miyake. Underdetermination, black boxes, and measurement. Philosophy of Science, 80(5):697–708, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/673729.

Teira:2013:CSE

[221] David Teira. A contractarian solution to the Experimenter’s regress. Philosophy of Science, 80(5):709–720, December 2013. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/673717.

Helgeson:2013:CSA

[222] Casey Helgeson. The confirmational significance of agreeing measurements. Philosophy of Science, 80(5):721–732, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/673924.

Sprenger:2013:TPN

[223] Jan Sprenger. Testing a precise null hypothesis: The case of Lindley’s paradox. Philosophy of Science, 80(5):733–744, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/673730.

Okasha:2013:EBU

[224] Samir Okasha. The evolution of Bayesian updating. Philosophy of Science, 80(5):745–757, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/674058.

Armendt:2013:PII

[225] Brad Armendt. Pragmatic interests and imprecise belief. Philosophy of Science, 80(5):758–768, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/673925.

Alexander:2013:PAS

[226] J. McKenzie Alexander. Preferential attachment and the search for successful theories. Philosophy of Science, 80(5):769–782, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/674080.

Wagner:2013:ERN

[227] Elliott Wagner. The explanatory relevance of Nash equilibrium: Onedimensional chaos in boundedly rational learning. Philosophy of Science, 80(5):783–795, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/673731.

Douglas:2013:VCV

[228] Heather Douglas. The value of cognitive values. Philosophy of Science, 80 (5):796–806, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
673716.

Elliott:2013:CAV

[229] Kevin C. Elliott and David Willmes. Cognitive attitudes and values in science. Philosophy of Science, 80(5):807–817, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/673719.

Steel:2013:AVI

[230] Daniel Steel. Acceptance, values, and inductive risk. Philosophy of Science, 80(5):818–828, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/673936.

Brown:2013:VSB

[231] Matthew J. Brown. Values in science beyond underdetermination and inductive risk. Philosophy of Science, 80(5):829–839, December 2013. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/673720.

Magnus:2013:WSK

[232] P. D. Magnus. What scientists know is not a function of what scientists know. Philosophy of Science, 80(5):840–849, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/673718.

Grune-Yanoff:2013:AMN

[233] Till Grüne-Yanoff. Appraising models nonrepresentationally. Philosophy of Science, 80(5):850–861, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/673893.

Goldsby:2013:SSL

[234] Michael Goldsby. The “structure” of the “strategy”: Looking at the Matthewson–Weisberg trade-off and its justificatory role for the multiplemodels approach. Philosophy of Science, 80(5):862–873, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/673728.

Knuuttila:2013:SMM

[235] Tarja Knuuttila and Andrea Loettgers. Synthetic modeling and mechanistic account: Material recombination and beyond. Philosophy of Science, 80(5):874–885, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/673965.

Frigg:2013:MIC

[236] Roman Frigg, Leonard A. Smith, and David A. Stainforth. The myopia of imperfect climate models: The case of UKCP09. Philosophy of Science, 80 (5):886–897, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
673892.

Wolff:2013:CLM

[237] Johanna Wolff. Are conservation laws metaphysically necessary? Philosophy of Science, 80(5):898–906, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/673715.

Theurer:2013:SCM

[238] Kari L. Theurer. Seventeenth-Century mechanism: An alternative framework for reductionism. Philosophy of Science, 80(5):907–918, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/673894.

Martin:2013:CID

[239] Joseph D. Martin. Is the contingentist/inevitabilist debate a matter of degrees? Philosophy of Science, 80(5):919–930, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/674003.

Sheredos:2013:WDB

[240] Benjamin Sheredos, Daniel Burnston, Adele Abrahamsen, and William Bechtel. Why do biologists use so many diagrams? Philosophy of Science, 80(5):931–944, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/674047.

Fagan:2013:SCU

[241] Melinda Bonnie Fagan. The stem cell uncertainty principle. Philosophy of Science, 80(5):945–957, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/674014.

Silberstein:2013:CLD

[242] Michael Silberstein and Anthony Chemero. Constraints on localization and decomposition as explanatory strategies in the biological sciences. Philosophy of Science, 80(5):958–970, December 2013. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/674533.

Barker:2013:DCA

[243] Matthew J. Barker and Joel D. Velasco. Deep conventionalism about evolutionary groups. Philosophy of Science, 80(5):971–982, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/673923.

Ramsey:2013:HNP

[244] Grant Ramsey. Human nature in a Post-essentialist world. Philosophy of Science, 80(5):983–993, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/673902.

Hochman:2013:DWN

[245] Adam Hochman. Do we need a device to acquire ethnic concepts? Philosophy of Science, 80(5):994–1005, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/673896.

Scarantino:2013:RFR

[246] Andrea Scarantino. Rethinking functional reference. Philosophy of Science, 80(5):1006–1018, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/673900.

Roche:2013:DTI

[247] Michael Roche. A difficulty for testing the inner sense theory of introspection. Philosophy of Science, 80(5):1019–1030, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/673967.

Boone:2013:OCS

[248] Worth Boone. Operationalizing consciousness: Subjective report and task performance. Philosophy of Science, 80(5):1031–1041, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/673966.

Eronen:2013:NLN

[249] Markus I. Eronen. No levels, no problems: Downward causation in neuroscience. Philosophy of Science, 80(5):1042–1052, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/673898.

Nardini:2013:BCS

[250] Cecilia Nardini and Jan Sprenger. Bias and conditioning in sequential medical trials. Philosophy of Science, 80(5):1053–1064, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/673732.

Tulodziecki:2013:SMS

[251] Dana Tulodziecki. Shattering the myth of semmelweis. Philosophy of Science, 80(5):1065–1075, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/673935.

Tamir:2013:GUG

[252] Michael Tamir. Geodesic universality in General Relativity. Philosophy of Science, 80(5):1076–1088, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/674057.

Muller:2013:GMT

[253] Thomas Müller. A generalized manifold topology for branching spacetimes. Philosophy of Science, 80(5):1089–1100, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/673895.

Brading:2013:PEH

[254] Katherine Brading. Presentism as an empirical hypothesis. Philosophy of Science, 80(5):1101–1111, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/673897.

Roberts:2013:WWD

[255] Bryan W. Roberts. When we do (and do not) have a classical arrow of time. Philosophy of Science, 80(5):1112–1124, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/674001.

Cuffaro:2013:DCP

[256] Michael E. Cuffaro and Wayne C. Myrvold. On the debate concerning the proper characterization of quantum dynamical evolution. Philosophy of Science, 80(5):1125–1136, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/673733.

Pashby:2013:DQO

[257] Thomas Pashby. Do quantum objects have temporal parts? Philosophy of Science, 80(5):1137–1147, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/673968.

Hemmo:2013:PZB

[258] Meir Hemmo and Orly Shenker. Probability zero in Bohm’s theory. Philosophy of Science, 80(5):1148–1158, December 2013. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/673891.

Goodwin:2013:QCO

[259] William Goodwin. Quantum chemistry and organic theory. Philosophy of Science, 80(5):1159–1169, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/673734.

Shech:2013:WPP

[260] Elay Shech. What is the paradox of phase transitions? Philosophy of Science, 80(5):1170–1181, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/674000.

Norton:2013:ETC

[261] John D. Norton. The end of the thermodynamics of computation: A no-go result. Philosophy of Science, 80(5):1182–1192, December 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/673714.

Elliott:2014:NVM

[262] Kevin C. Elliott and Daniel J. McKaughan. Nonepistemic values and the multiple goals of science. Philosophy of Science, 81(1):1–21, January 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/674345.

Sarkar:2014:DIP

[263] Sahotra Sarkar. Does “information” provide a compelling framework for a theory of natural selection? Grounds for caution. Philosophy of Science, 81(1):22–30, January 2014. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
674204.

Frigg:2014:LDA

[264] Roman Frigg, Seamus Bradley, Hailiang Du, and Leonard A. Smith. Laplace’s demon and the adventures of his apprentices. Philosophy of Science, 81(1):31–59, January 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/674416.

Russell:2014:WTC

[265] Jeffrey Sanford Russell. On where things could be. Philosophy of Science, 81(1):60–80, January 2014. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
674344.

Rinard:2014:NBS

[266] Susanna Rinard. A new Bayesian solution to the paradox of the ravens. Philosophy of Science, 81(1):81–100, January 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.

jstor.org/stable/10.1086/674202.

vanEnk:2014:BMC

[267] Steven J. van Enk. Bayesian measures of confirmation from scoring rules. Philosophy of Science, 81(1):101–113, January 2014. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/674205.

Roche:2014:DC

[268] William Roche and Tomoji Shogenji. Dwindling confirmation. Philosophy of Science, 81(1):114–137, January 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/674346.

Gebharter:2014:FFR

[269] Alexander Gebharter. A formal framework for representing mechanisms? Philosophy of Science, 81(1):138–153, January 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/674206.

Michael:2014:WIA

[270] John Michael and Francesca Fardo. What (if anything) is shared in pain empathy? A critical discussion of de Vignemont and Jacob’s theory of the neural substrate of pain empathy. Philosophy of Science, 81(1):154– 160, January 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/674203. See

[121].

Isaac:2014:BRP

[271] Alistair M. C. Isaac. Book review: Paul M. Churchland, Plato’s Camera: How the Physical Brain Captures a Landscape of Abstract Universals. Philosophy of Science, 81(1):161–165, January 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/674366.

Wilholt:2014:BRP

[272] Torsten Wilholt. Book review: Philip Kitcher, Science in a Democratic Society. Philosophy of Science, 81(1):165–171, January 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/674367.

Baker:2014:BRF

[273] David John Baker. Book review: Frank Arntzenius, Space, Time, and Stuff. Philosophy of Science, 81(1):171–174, January 2014. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/674433.

Sterner:2014:PVB

[274] Beckett Sterner. The practical value of biological information for research. Philosophy of Science, 81(2):175–194, April 2014. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/675679.

Clarke:2014:NCT

[275] Christopher Clarke. Neuroeconomics and confirmation theory. Philosophy of Science, 81(2):195–215, April 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/675669.

Robert:2014:JLP

[276] Christian P. Robert. On the Jeffreys–Lindley paradox. Philosophy of Science, 81(2):216–232, April 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/675729.

Weatherall:2014:GC

[277] James Owen Weatherall and John Byron Manchak. The geometry of conventionality. Philosophy of Science, 81(2):233–247, April 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/675680.

Esfeld:2014:GFT

[278] Michael Esfeld and Nicolas Gisin. The GRW flash theory: A relativistic quantum ontology of matter in space–time? Philosophy of Science, 81 (2):248–264, April 2014. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
675730.

Smith:2014:DBA

[279] Ryan Smith. Do brains have an arrow of time? Philosophy of Science, 81(2):265–275, April 2014. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
675644.

vanFraassen:2014:OTG

[280] Bas C. van Fraassen. One or two gentle remarks about Hans Halvorson’s critique of the semantic view. Philosophy of Science, 81(2):276–283, April 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/675645.

Ruse:2014:BRP

[281] Michael Ruse. Book review: Paul Hoyningen-Huene, Systematicity: The Nature of Science. Philosophy of Science, 81(2):284–288, April 2014. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/676562.

Sklar:2014:BRT

[282] Lawrence Sklar. Book review: Tim Maudlin, Philosophy of Physics: Space and Time. Philosophy of Science, 81(2):288–292, April 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/676563.

Odenbaugh:2014:BRS

[283] Jay Odenbaugh. Book review: Sahotra Sarkar, Environmental Philosophy: From Theory to Practice. Philosophy of Science, 81(2):292–296, April 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/676561.

Anonymous:2014:RBS

[284] Anonymous. Referees for Philosophy of Science. Philosophy of Science, 81(2):297–301, April 2014. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
676320.

Skyrms:2014:GIS

[285] Brian Skyrms. Grades of inductive skepticism. Philosophy of Science, 81(3):303–312, July 2014. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
676637.

Jantzen:2014:PVT

[286] Benjamin C. Jantzen. Piecewise versus total support: How to deal with background information in likelihood arguments. Philosophy of Science, 81(3):313–331, July 2014. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
676536.

Jansson:2014:CTE

[287] Lina Jansson. Causal theories of explanation and the challenge of explanatory disagreement. Philosophy of Science, 81(3):332–348, July 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/676687.

Batterman:2014:MME

[288] Robert W. Batterman and Collin C. Rice. Minimal model explanations. Philosophy of Science, 81(3):349–376, July 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/676677. See reply [366].

Peters:2014:WES

[289] Dean Peters. What elements of successful scientific theories are the correct targets for “selective” scientific realism? Philosophy of Science, 81(3):377– 397, July 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/676537.

Santana:2014:ACS

[290] Carlos Santana. Ambiguity in cooperative signaling. Philosophy of Science, 81(3):398–422, July 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/676652.

Stan:2014:UKN

[291] Marius Stan. Unity for Kant’s natural philosophy. Philosophy of Science, 81(3):423–443, July 2014. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
676538.

DeLanghe:2014:UMD

[292] Rogier De Langhe. A unified model of the division of cognitive labor. Philosophy of Science, 81(3):444–459, July 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/676670.

Walsh:2014:CRC

[293] Sean Walsh, Eleanor Knox, and Adam Caulton. Critical review: Christopher Pincock, Mathematics and Scientific Representation. Philosophy of Science, 81(3):460–469, July 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/676560.

Falk:2014:BRP

[294] Raphael Falk. Book review: Paul Griffiths and Karola Stotz, Genetics and Philosophy: An Introduction. Philosophy of Science, 81(3):470–475, July 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/677259.

Hillier:2014:BRG

[295] Sam Hillier. Book review: Greg Frost-Arnold, Carnap, Tarski, and Quine at Harvard. Philosophy of Science, 81(3):476–480, July 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/677258.

Dunlop:2014:BRJ

[296] Katherine Dunlop. Book review: Jeremy Gray, Henri Poincaré: A Scientific Biography. Philosophy of Science, 81(3):481–486, July 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/677256.

Klein:2014:BRA

[297] Colin Klein. Book review: Alcino J. Silva, Anthony Landreth, and John Bickle, Engineering the Next Revolution in Neuroscience. Philosophy of Science, 81(3):486–489, July 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/677257.

Bradley:2014:ACJ

[298] Richard Bradley, Franz Dietrich, and Christian List. Aggregating causal judgments. Philosophy of Science, 81(4):491–515, October 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/678044.

Parke:2014:ESE

[299] Emily C. Parke. Experiments, simulations, and epistemic privilege. Philosophy of Science, 81(4):516–536, October 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/677956.

Huisman:2014:IUC

[300] Leendert Huisman. On indeterminate updating of credences. Philosophy of Science, 81(4):537–557, October 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/677955.

Sober:2014:TKE

[301] Elliott Sober and Mike Steel. Time and knowability in evolutionary processes. Philosophy of Science, 81(4):558–579, October 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/677954.

Artiga:2014:MTF

[302] Marc Artiga. The modal theory of function is not about functions. Philosophy of Science, 81(4):580–591, October 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/677952.

Lamm:2014:IC

[303] Ehud Lamm and Ohad Kammar. Inferring coevolution. Philosophy of Science, 81(4):592–611, October 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/678045.

Booth:2014:PIH

[304] Austin Booth. Populations and individuals in heterokaryotic fungi: A multilevel perspective. Philosophy of Science, 81(4):612–632, October 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/677953.

Bursten:2014:MEN

[305] Julia R. Bursten. Microstructure without essentialism: A new perspective on chemical classification. Philosophy of Science, 81(4):633–653, October 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/678043.

Ehrlich:2014:EHA

[306] Philip Ehrlich. An essay in honor of Adolf Grünbaum’s ninetieth birthday: A reexamination of Zeno’s paradox of extension. Philosophy of Science, 81(4):654–675, October 2014. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
677978.

Driscoll:2014:BRB

[307] Catherine Driscoll. Book review: Studying Human Behavior: How Scientists Investigate Aggression and Sexuality. Philosophy of Science, 81 (4):676–680, October 2014. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
678555.

Cuffaro:2014:BRC

[308] Michael E. Cuffaro. Book review: Christopher G. Timpson, Quantum Information Theory and the Foundations of Quantum Mechanics. Philosophy of Science, 81(4):681–684, October 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/678554.

Harper:2014:BRP

[309] William Harper. Book review: Peter Achinstein, Evidence and Method: Scientific Strategies of Isaac Newton and James Clerk Maxwell. Philosophy of Science, 81(4):684–687, October 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/678556.

Woody:2014:P

[310] Andrea I. Woody. Preface. Philosophy of Science, 81(5):689–690, December 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/678333.

Woodward:2014:FAC

[311] James Woodward. A functional account of causation; or, a defense of the legitimacy of causal thinking by reference to the only standard that matters — usefulness (as opposed to metaphysics or agreement with intuitive judgment). Philosophy of Science, 81(5):691–713, December 2014. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/678313.

Ereshefsky:2014:SHP

[312] Marc Ereshefsky. Species, historicity, and path dependency. Philosophy of Science, 81(5):714–726, December 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/677202.

Charbonneau:2014:PR

[313] Mathieu Charbonneau. Populations without reproduction. Philosophy of Science, 81(5):727–740, December 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/677203.

Millstein:2014:HCP

[314] Roberta L. Millstein. How the concept of population resolves concepts of environment. Philosophy of Science, 81(5):741–755, December 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/677220.

Weber:2014:EMB

[315] Marcel Weber. Experimental modeling in biology: In vivo representation and stand-ins as modeling strategies. Philosophy of Science, 81(5): 756–769, December 2014. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
678257.

Justus:2014:MIE

[316] James Justus. Methodological individualism in ecology. Philosophy of Science, 81(5):770–784, December 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/677404.

Weisberg:2014:UEP

[317] Michael Weisberg. Understanding the emergence of population behavior in individual-based models. Philosophy of Science, 81(5):785–797, December 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/677405.

Nanay:2014:TE

[318] Bence Nanay. Teleosemantics without etiology. Philosophy of Science, 81 (5):798–810, December 2014. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
677684.

OMalley:2014:MRS

[319] Maureen A. O’Malley, Ingo Brigandt, Alan C. Love, John W. Crawford, Jack A. Gilbert, Rob Knight, Sandra D. Mitchell, and Forest Rohwer. Multilevel research strategies and biological systems. Philosophy of Science, 81(5):811–828, December 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/677889.

Barrett:2014:RFE

[320] Jeffrey A. Barrett. Rule-following and the evolution of basic concepts. Philosophy of Science, 81(5):829–839, December 2014. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/678241.

OConnor:2014:EPC

[321] Cailin O’Connor. Evolving perceptual categories. Philosophy of Science, 81(5):840–851, December 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/677885.

Smead:2014:DEP

[322] Rory Smead. Deception and the evolution of plasticity. Philosophy of Science, 81(5):852–865, December 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/677685.

Godfrey-Smith:2014:SRS

[323] Peter Godfrey-Smith. Sender-receiver systems within and between organisms. Philosophy of Science, 81(5):866–878, December 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/677686.

Calcott:2014:CRI

[324] Brett Calcott. The creation and reuse of information in gene regulatory networks. Philosophy of Science, 81(5):879–890, December 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/677687.

Cao:2014:SBS

[325] Rosa Cao. Signaling in the brain: In search of functional units. Philosophy of Science, 81(5):891–901, December 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/677688.

Shea:2014:NSP

[326] Nicholas Shea. Neural signaling of probabilistic vectors. Philosophy of Science, 81(5):902–913, December 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/678354.

Chirimuuta:2014:PME

[327] M. Chirimuuta. Psychophysical methods and the evasion of introspection. Philosophy of Science, 81(5):914–926, December 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/677890.

Feest:2014:PEF

[328] Uljana Feest. Phenomenal experiences, first-person methods, and the artificiality of experimental data. Philosophy of Science, 81(5):927–939, December 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/677689.

Hatfield:2014:PEP

[329] Gary Hatfield. Psychological experiments and phenomenal experience in size and shape constancy. Philosophy of Science, 81(5):940–953, December 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/677891.

Schwitzgebel:2014:PKI

[330] Eric Schwitzgebel. The problem of known illusion and the resemblance of experience to reality. Philosophy of Science, 81(5):954–960, December 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/677690.

Aktunc:2014:STN

[331] M. Emrah Aktunc. Severe tests in neuroimaging: What we can learn and how we can learn it. Philosophy of Science, 81(5):961–973, December 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/677691.

Klein:2014:BRW

[332] Colin Klein. The brain at rest: What it is doing and why that matters. Philosophy of Science, 81(5):974–985, December 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/677692.

Schwartz:2014:STR

[333] Peter H. Schwartz. Small tumors as risk factors not disease. Philosophy of Science, 81(5):986–998, December 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/678280.

Ankeny:2014:ORC

[334] Rachel A. Ankeny. The overlooked role of cases in casual attribution in medicine. Philosophy of Science, 81(5):999–1011, December 2014. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/677693.

Morgan:2014:RKG

[335] Mary S. Morgan. Resituating knowledge: Generic strategies and case studies. Philosophy of Science, 81(5):1012–1024, December 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/677888.

Spencer:2014:RSR

[336] Quayshawn Spencer. A radical solution to the race problem. Philosophy of Science, 81(5):1025–1038, December 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/677694.

Kaplan:2014:RAC

[337] Jonathan Michael Kaplan and Rasmus Grønfeldt Winther. Realism, antirealism, and conventionalism about race. Philosophy of Science, 81(5): 1039–1052, December 2014. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
678314.

Kopec:2014:CCC

[338] Matthew Kopec. Clines, clusters, and clades in the race debate. Philosophy of Science, 81(5):1053–1065, December 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/677695.

Manchak:2014:STS

[339] John Byron Manchak. On space-time singularities, holes, and extensions. Philosophy of Science, 81(5):1066–1076, December 2014. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/677696.

Weatherall:2014:WSG

[340] James Owen Weatherall. What is a singularity in geometrized Newtonian gravitation? Philosophy of Science, 81(5):1077–1089, December 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/678239.

McKenzie:2014:FS

[341] Kerry McKenzie. On the fundamentality of symmetries. Philosophy of Science, 81(5):1090–1102, December 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/677697.

Humphreys:2014:ECS

[342] Paul Humphreys. Explanation as condition satisfaction. Philosophy of Science, 81(5):1103–1116, December 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/677698.

Kuhlmann:2014:EFM

[343] Meinard Kuhlmann. Explaining financial markets in terms of complex systems. Philosophy of Science, 81(5):1117–1130, December 2014. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/677699.

Bedau:2014:TBM

[344] Mark A. Bedau. Testing bottom-up models of complex citation networks. Philosophy of Science, 81(5):1131–1143, December 2014. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/677700.

Morrison:2014:CSR

[345] Margaret Morrison. Complex systems and renormalization group explanations. Philosophy of Science, 81(5):1144–1156, December 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/677904.

Reutlinger:2014:WTU

[346] Alexander Reutlinger. Why is there universal macrobehavior? Renormalization group explanation as noncausal explanation. Philosophy of Science, 81(5):1157–1170, December 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/677887.

Lutz:2014:EAR

[347] Sebastian Lutz. Empirical adequacy in the received view. Philosophy of Science, 81(5):1171–1183, December 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/677886.

Hardcastle:2014:NET

[348] Gary Hardcastle and Matthew H. Slater. A novel exercise for teaching the philosophy of science. Philosophy of Science, 81(5):1184–1196, December 2014. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/678240.

Haufe:2015:GL

[349] Chris Haufe. Gould’s Laws. Philosophy of Science, 82(1):1–20, January 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/678979.

Helgeson:2015:TNA

[350] Casey Helgeson. There is no asymmetry of identity assumptions in the debate over selection and individuals. Philosophy of Science, 82(1):21–31, January 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/679114.

Ross:2015:DME

[351] Lauren N. Ross. Dynamical models and explanation in neuroscience. Philosophy of Science, 82(1):32–54, January 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/679038.

Hey:2015:RDE

[352] Spencer Phillips Hey. Robust and discordant evidence: Methodological lessons from clinical research. Philosophy of Science, 82(1):55–75, January 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/678978.

Dawid:2015:HDL

[353] Richard Dawid. Higgs discovery and the look elsewhere effect. Philosophy of Science, 82(1):76–96, January 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/679179.

Cohen:2015:SSM

[354] Michael P. Cohen. On Schupbach and Sprenger’s measures of explanatory power. Philosophy of Science, 82(1):97–109, January 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/678980.

OConnor:2015:AKG

[355] Cailin O’Connor. Ambiguity is kinda good sometimes. Philosophy of Science, 82(1):110–121, January 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/679180.

Crawford:2015:SF

[356] David Robert Crawford. Sarkar on Frank. Philosophy of Science, 82 (1):122–128, January 2015. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
679035.

Bacciagaluppi:2015:BRB

[357] Guido Bacciagaluppi and Jenann Ismael. Book review: The Emergent Multiverse: Quantum Theory according to the Everett Interpretation. Philosophy of Science, 82(1):129–148, January 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/679037.

Maudlin:2015:BRP

[358] Tim Maudlin. Book review: L. A. Paul and Ned Hall, Causation: A User’s Guide. Philosophy of Science, 82(1):149–152, January 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/678957.

Balcells:2015:BRA

[359] Maria Balcells. Book review: Adrian Bardon, A Brief History of the Philosophy of Time. Philosophy of Science, 82(1):152–155, January 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/679036.

Rolin:2015:VSC

[360] Kristina Rolin. Values in science: The case of scientific collaboration. Philosophy of Science, 82(2):157–177, April 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/680522.

Matthewson:2015:DPD

[361] John Matthewson. Defining paradigm Darwinian populations. Philosophy of Science, 82(2):178–197, April 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/680665.

Ratti:2015:BDB

[362] Emanuele Ratti. Big data biology: Between eliminative inferences and exploratory experiments. Philosophy of Science, 82(2):198–218, April 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/680332.

Grim:2015:GGM

[363] Patrick Grim, Daniel J. Singer, Christopher Reade, and Steven Fisher. Germs, genes, and memes: Function and fitness dynamics on information networks. Philosophy of Science, 82(2):219–243, April 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/680486.

Ludwig:2015:ANM

[364] David Ludwig. Against the new metaphysics of race. Philosophy of Science, 82(2):244–265, April 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/680487.

Sebens:2015:QMC

[365] Charles T. Sebens. Quantum mechanics as classical physics. Philosophy of Science, 82(2):266–291, April 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/680190.

Lange:2015:MME

[366] Marc Lange. On “minimal model explanations”: A reply to Batterman and Rice. Philosophy of Science, 82(2):292–305, April 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/680488. See [288].

Klee:2015:BRB

[367] Robert Klee. Book review: Seeing Things: The Philosophy of Reliable Observation. Philosophy of Science, 82(2):306–312, April 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/680387.

Schwarz:2015:BRB

[368] Wolfgang Schwarz. Book review: Tychomancy: Inferring Probability from Causal Structure. Philosophy of Science, 82(2):313–320, April 2015. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/680490.

Levy:2015:BRB

[369] Arnon Levy. Book review: In Search of Mechanisms. Philosophy of Science, 82(2):321–325, April 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/680385.

Butterfield:2015:BRA

[370] Jeremy Butterfield. Book review: Anthony Duncan, The Conceptual Framework of Quantum Field Theory. Philosophy of Science, 82(2):326– 330, April 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/680386.

Blanchard:2015:BRD

[371] Thomas Blanchard. Book review: Douglas Kutach, Causation and Its Basis in Fundamental Physics. Philosophy of Science, 82(2):330–333, April 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/680489.

Anonymous:2015:E

[372] Anonymous. Erratum. Philosophy of Science, 82(2):334, April 2015. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/680346.

Anonymous:2015:RBS

[373] Anonymous. Referees for Philosophy of Science. Philosophy of Science, 82(2):335–339, April 2015. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
680497.

Reiss:2015:PTE

[374] Julian Reiss. A pragmatist theory of evidence. Philosophy of Science, 82(3):341–362, July 2015. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
681643.

Gouvea:2015:EEF

[375] Devin Y. Gouvêa. Explanation and the evolutionary first law(s). Philosophy of Science, 82(3):363–382, July 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/681603.

Sprenger:2015:NSP

[376] Jan Sprenger. A novel solution to the problem of old evidence. Philosophy of Science, 82(3):383–401, July 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/681767.

Sterelny:2015:OER

[377] Kim Sterelny. Optimizing engines: Rational choice in the Neolithic? Philosophy of Science, 82(3):402–423, July 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/681602.

Alexander:2015:ELO

[378] Jason McKenzie Alexander, Johannes Himmelreich, and Christopher Thompson. Epistemic landscapes, optimal search, and the division of cognitive labor. Philosophy of Science, 82(3):424–453, July 2015. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/681766.

Thoma:2015:EDL

[379] Johanna Thoma. The epistemic division of labor revisited. Philosophy of Science, 82(3):454–472, July 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/681768.

Halina:2015:TNS

[380] Marta Halina. There is no special problem of mindreading in nonhuman animals. Philosophy of Science, 82(3):473–490, July 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/681627.

Krueger:2015:THM

[381] James Krueger. Theoretical health and medical practice. Philosophy of Science, 82(3):491–508, July 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/681628.

Dammann:2015:BRE

[382] Olaf Dammann. Book review: Epidemiological explanations: Philosophy of Epidemiology, by Alex Broadbent. Philosophy of Science, 82(3):509– 519, July 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/681769.

Bhakthavatsalam:2015:BRS

[383] Sindhuja Bhakthavatsalam. Book review: Steven J. Dick, Discovery and Classification in Astronomy — Controversy and Consensus. Philosophy of Science, 82(3):520–524, July 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/681770.

Elliott:2015:BRD

[384] Kevin Elliott. Book review: Daniel Steel, Philosophy and the Precautionary Principle: Science, Evidence, and Environmental Policy. Philosophy of Science, 82(3):524–527, July 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/681642.

Griffiths:2015:MCS

[385] Paul E. Griffiths, Arnaud Pocheville, Brett Calcott, Karola Stotz, Hyunju Kim, and Rob Knight. Measuring causal specificity. Philosophy of Science, 82(4):529–555, October 2015. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
682914.

Murray-Watters:2015:WGI

[386] Alexander Murray-Watters and Clark Glymour. What is going on inside the arrows? Discovering the hidden springs in causal models. Philosophy of Science, 82(4):556–586, October 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/682962.

Huttegger:2015:BCT

[387] Simon M. Huttegger. Bayesian convergence to the truth and the metaphysics of possible worlds. Philosophy of Science, 82(4):587–601, October 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/682941.

Stefansson:2015:HVC

[388] H. Orri Stefánsson and Richard Bradley. How valuable are chances? Philosophy of Science, 82(4):602–625, October 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/682915.

Caie:2015:CIC

[389] Michael Caie. Credence in the image of chance. Philosophy of Science, 82 (4):626–648, October 2015. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
682963.

Bruner:2015:DIT

[390] Justin P. Bruner. Disclosure and information transfer in signaling games. Philosophy of Science, 82(4):649–666, October 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.

jstor.org/stable/10.1086/683016.

Boyer-Kassem:2015:SCD

[391] Thomas Boyer-Kassem and Cyrille Imbert. Scientific collaboration: Do two heads need to be more than twice better than one? Philosophy of Science, 82(4):667–688, October 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/682940.

Hoover:2015:REI

[392] Kevin D. Hoover. Reductionism in economics: Intentionality and eschatological justification in the microfoundations of macroeconomics. Philosophy of Science, 82(4):689–711, October 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/682917.

Hartmann:2015:NGS

[393] Stephan Hartmann and Branden Fitelson. A new Garber-style solution to the problem of old evidence. Philosophy of Science, 82(4):712–717, October 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/682916.

Smeenk:2015:BRB

[394] Chris Smeenk. Book review: Kant’s Construction of Nature: A Reading of the Metaphysical Foundations of Natural Science. Philosophy of Science, 82(4):718–726, October 2015. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
682991.

Pronskikh:2015:BRA

[395] Vitaly Pronskikh. Book review: Allan Franklin, Shifting Standards: Experiments in Particle Physics in the Twentieth Century. Philosophy of Science, 82(4):727–730, October 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/682990.

OConnor:2015:BRP

[396] Cailin O’Connor. Book review: Peter Godfrey-Smith, Philosophy of Biology. Philosophy of Science, 82(4):731–733, October 2015. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/682992.

Smeenk:2015:P

[397] Chris Smeenk. Preface. Philosophy of Science, 82(5):735–736, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/684164.

Cuffaro:2015:HPE

[398] Michael E. Cuffaro. How-possibly explanations in (quantum) computer science. Philosophy of Science, 82(5):737–748, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/683243.

Nyrup:2015:HER

[399] Rune Nyrup. How explanatory reasoning justifies pursuit: A Peircean view of IBE. Philosophy of Science, 82(5):749–760, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/683262.

Love:2015:ICM

[400] Alan C. Love and Marco J. Nathan. The idealization of causation in mechanistic explanation. Philosophy of Science, 82(5):761–774, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/683263.

Baetu:2015:CME

[401] Tudor M. Baetu. The completeness of mechanistic explanations. Philosophy of Science, 82(5):775–786, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/683279.

Heilmann:2015:NIR

[402] Conrad Heilmann. A new interpretation of the representational theory of measurement. Philosophy of Science, 82(5):787–797, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/683280.

Fumagalli:2015:NLM

[403] Roberto Fumagalli. No learning from minimal models. Philosophy of Science, 82(5):798–809, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/683281.

Leonelli:2015:WCS

[404] Sabina Leonelli. What counts as scientific data? A relational framework. Philosophy of Science, 82(5):810–821, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/684083.

Miyake:2015:RMU

[405] Teru Miyake. Reference models: Using models to turn data into evidence. Philosophy of Science, 82(5):822–832, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/683322.

Perry:2015:PEQ

[406] Zee R. Perry. Properly extensive quantities. Philosophy of Science, 82 (5):833–844, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
683323.

Robus:2015:DSL

[407] Olin M. Robus. Does science license metaphysics? Philosophy of Science, 82(5):845–855, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/684023.

Sample:2015:SUA

[408] Matthew S. Sample. Stanford’s unconceived alternatives from the perspective of epistemic obligations. Philosophy of Science, 82(5):856–866, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/683324.

Stanford:2015:CUS

[409] P. Kyle Stanford. Catastrophism, uniformitarianism, and a scientific realism debate that makes a difference. Philosophy of Science, 82(5):867–878, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/683325.

Benetreau-Dupin:2015:BCP

[410] Yann Benétreau-Dupin. Blurring out cosmic puzzles. Philosophy of Science, 82(5):879–891, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/683326.

Vassend:2015:CMS

[411] Olav B. Vassend. Confirmation measures and sensitivity. Philosophy of Science, 82(5):892–904, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/683327.

Pietsch:2015:ATL

[412] Wolfgang Pietsch. Aspects of theory-ladenness in data-intensive science. Philosophy of Science, 82(5):905–916, December 2015. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/683328.

Malinsky:2015:HTD

[413] Daniel Malinsky. Hypothesis testing, “Dutch book” arguments, and risk. Philosophy of Science, 82(5):917–929, December 2015. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/683341.

Zhang:2015:LCF

[414] Jiji Zhang and Kun Zhang. Likelihood and consilience: On Forster’s counterexamples to the likelihood theory of evidence. Philosophy of Science, 82 (5):930–940, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
683342.

Fillion:2015:NMC

[415] Nicolas Fillion and Sorin Bangu. Numerical methods, complexity, and epistemic hierarchies. Philosophy of Science, 82(5):941–955, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/683343.

Holman:2015:PIB

[416] Bennett Holman and Justin P. Bruner. The problem of intransigently biased agents. Philosophy of Science, 82(5):956–968, December 2015. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/683344.

Alexander:2015:CTR

[417] J. McKenzie Alexander. Cheap talk, reinforcement learning, and the emergence of cooperation. Philosophy of Science, 82(5):969–982, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/684197.

Abrams:2015:CMR

[418] Marshall Abrams. Coherence, Muller’s ratchet, and the maintenance of culture. Philosophy of Science, 82(5):983–996, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/683434.

Huttegger:2015:HPA

[419] Simon M. Huttegger, Justin P. Bruner, and Kevin J. S. Zollman. The handicap principle is an artifact. Philosophy of Science, 82(5):997–1009, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/683435.

McCaffrey:2015:BHF

[420] Joseph B. McCaffrey. The brain’s heterogeneous functional landscape. Philosophy of Science, 82(5):1010–1022, December 2015. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/683436.

Zautra:2015:EIE

[421] Nicholas Zautra. Embodiment, interaction, and experience: Toward a comprehensive model in addiction science. Philosophy of Science, 82(5): 1023–1034, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
683437.

Povich:2015:MMB

[422] Mark Povich. Mechanisms and model-based functional magnetic resonance imaging. Philosophy of Science, 82(5):1035–1046, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/683438.

Tabb:2015:PPA

[423] Kathryn Tabb. Psychiatric progress and the assumption of diagnostic discrimination. Philosophy of Science, 82(5):1047–1058, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/683439.

Mikhalevich:2015:EAM

[424] Irina Mikhalevich. Experiment and animal minds: Why the choice of the null hypothesis matters. Philosophy of Science, 82(5):1059–1069, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/683440.

ONeill:2015:WCM

[425] Elizabeth O’Neill. Which causes of moral beliefs matter? Philosophy of Science, 82(5):1070–1080, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/683441.

Pence:2015:OFB

[426] Charles H. Pence and Grant Ramsey. Is organismic fitness at the basis of evolutionary theory? Philosophy of Science, 82(5):1081–1091, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/683442.

Kovaka:2015:BIS

[427] Karen Kovaka. Biological individuality and scientific practice. Philosophy of Science, 82(5):1092–1103, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/683443.

Linquist:2015:ALC

[428] Stefan Linquist. Against Lawton’s contingency thesis; or, why the reported demise of community ecology is greatly exaggerated. Philosophy of Science, 82(5):1104–1115, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/684024.

Matthews:2015:EMP

[429] Lucas J. Matthews. Embedded mechanisms and phylogenetics. Philosophy of Science, 82(5):1116–1126, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/683444.

Kincaid:2015:OEM

[430] Harold Kincaid. Open empirical and methodological issues in the individualism–holism debate. Philosophy of Science, 82(5):1127–1138, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/683445.

Skillings:2015:MEB

[431] Derek John Skillings. Mechanistic explanation of biological processes. Philosophy of Science, 82(5):1139–1151, December 2015. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/683446.

Stevens:2015:DAP

[432] Syman Stevens. The dynamical approach as practical geometry. Philosophy of Science, 82(5):1152–1162, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/683447.

Huggett:2015:DGR

[433] Nick Huggett and Tiziana Vistarini. Deriving general relativity from string theory. Philosophy of Science, 82(5):1163–1174, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/683448.

Rosaler:2015:BBT

[434] Joshua Rosaler. Is de Broglie–Bohm theory specially equipped to recover classical behavior? Philosophy of Science, 82(5):1175–1187, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/683468.

Norton:2015:WDR

[435] Joshua Norton. Weak discernibility and relations between quanta. Philosophy of Science, 82(5):1188–1199, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/683647.

Kao:2015:UQH

[436] Molly Kao. Unification and the quantum hypothesis in 1900–1913. Philosophy of Science, 82(5):1200–1210, December 2015. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/683648.

Li:2015:CGR

[437] Bihui Li. Coarse-graining as a route to microscopic physics: The renormalization group in quantum field theory. Philosophy of Science, 82(5): 1211–1223, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
684085.

Werndl:2015:RBE

[438] Charlotte Werndl and Roman Frigg. Rethinking Boltzmannian equilibrium. Philosophy of Science, 82(5):1224–1235, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/683649.

Mayo-Wilson:2015:SC

[439] Conor Mayo-Wilson. Structural chaos. Philosophy of Science, 82(5): 1236–1247, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
684086.

Lombardi:2015:PVA

[440] Olimpia Lombardi, Sebastian Fortin, and Leonardo Vanni. A pluralist view about information. Philosophy of Science, 82(5):1248–1259, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/10.1086/683650.

Northcott:2015:OPE

[441] Robert Northcott. Opinion polling and election predictions. Philosophy of Science, 82(5):1260–1271, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/683651.

Lee:2015:CBP

[442] Carole J. Lee. Commensuration bias in peer review. Philosophy of Science, 82(5):1272–1283, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/683652.

Keren:2015:SIC

[443] Arnon Keren. Science and informed, counterfactual, democratic consent. Philosophy of Science, 82(5):1284–1295, December 2015. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/683653.

Runhardt:2015:ECM

[444] Rosa W. Runhardt. Evidence for causal mechanisms in social science: Recommendations from Woodward’s manipulability theory of causation. Philosophy of Science, 82(5):1296–1307, December 2015. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/683679.

Marcellesi:2015:EVT

[445] Alexandre Marcellesi. External validity: Is there still a problem? Philosophy of Science, 82(5):1308–1317, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/684084.

Hey:2015:WTT

[446] Spencer Phillips Hey. What theories are tested in clinical trials? Philosophy of Science, 82(5):1318–1329, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/683816.

Holman:2015:WMS

[447] Bennett Holman. Why most sugar pills are not placebos. Philosophy of Science, 82(5):1330–1343, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/683817.

Bradley:2016:CFE

[448] Seamus Bradley and Katie Steele. Can free evidence be bad? Value of information for the imprecise probabilist. Philosophy of Science, 83(1):1– 28, January 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/684184.

Vanderschraaf:2016:WDS

[449] Peter Vanderschraaf. In a weakly dominated strategy is strength: Evolution of optimality in stag hunt augmented with a punishment option. Philosophy of Science, 83(1):29–59, January 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/684166.

Bright:2016:CII

[450] Liam Kofi Bright, Daniel Malinsky, and Morgan Thompson. Causally interpreting intersectionality theory. Philosophy of Science, 83(1):60–81, January 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/684173.

Weatherall:2016:MHN

[451] James Owen Weatherall. Maxwell–Huygens, Newton–Cartan, and Saunders–Knox space–times. Philosophy of Science, 83(1):82–92, January 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/684080.

Teh:2016:GGU

[452] Nicholas J. Teh. Galileo’s gauge: Understanding the empirical significance of gauge symmetry. Philosophy of Science, 83(1):93–118, January 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/684196.

Thompson:2016:CTT

[453] J. Robert Thompson. Causing trouble: Theories of reference and theory of mind. Philosophy of Science, 83(1):119–130, January 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/684081.

Fresco:2016:IVI

[454] Nir Fresco and Michaelis Michael. Information and veridicality: Information processing and the Bar-Hillel/Carnap Paradox. Philosophy of Science, 83(1):131–151, January 2016. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
684165.

Serban:2016:BRW

[455] Maria Serban and Sara Green. Book review: Why the Small Things in Life Matter: Philosophy of Biology from the Microbial Perspective. Philosophy of Science, 83(1):152–158, January 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/684100.

Stefansson:2016:BRR

[456] H. Orri Stefánsson. Book review: Review of Evidence, Decision and Causality. Philosophy of Science, 83(1):159–169, January 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/684183.

Anonymous:2016:E

[457] Anonymous. Erratum. Philosophy of Science, 83(1):170, January 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/684247.

Nguyen:2016:PEB

[458] James Nguyen. On the pragmatic equivalence between representing data and phenomena. Philosophy of Science, 83(2):171–191, April 2016. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Clarke:2016:PPM

[459] Christopher Clarke. Preferences and positivist methodology in economics. Philosophy of Science, 83(2):192–212, April 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Halpern:2016:SCC

[460] Joseph Y. Halpern. Sufficient conditions for causality to be transitive. Philosophy of Science, 83(2):213–226, April 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Kuorikoski:2016:EDT

[461] Jaakko Kuorikoski and Caterina Marchionni. Evidential diversity and the triangulation of phenomena. Philosophy of Science, 83(2):227–247, April 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Demarest:2016:UOC

[462] Heather Demarest. The universe had one chance. Philosophy of Science, 83(2):248–264, April 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Manchak:2016:ETH

[463] John Byron Manchak. Epistemic “holes”’ in space–time. Philosophy of Science, 83(2):265–276, April 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Stan:2016:HIS

[464] Marius Stan. Huygens on inertial structure and relativity. Philosophy of Science, 83(2):277–298, April 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anonymous:2016:RPS

[465] Anonymous. Referees for Philosophy of Science. Philosophy of Science, 83(2):299–303, April 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anonymous:2016:Ma

[466] Anonymous. Masthead. Philosophy of Science, 83(3):??, July 2016. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Elga:2016:BH

[467] Adam Elga. Bayesian humility. Philosophy of Science, 83(3):305–323, July 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Cisewski:2016:SBC

[468] Jessi Cisewski, Joseph B. Kadane, Mark J. Schervish, Teddy Seidenfeld, and Rafael Stern. Sleeping Beauty’s credences. Philosophy of Science, 83(3):324–347, July 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Icard:2016:PCC

[469] Thomas F. Icard III. Pragmatic considerations on comparative probability. Philosophy of Science, 83(3):348–370, July 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Martinez:2016:CIS

[470] Manolo Martínez and Peter Godfrey-Smith. Common interest and signaling games: a dynamic analysis. Philosophy of Science, 83(3):371–392, July 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Luczak:2016:HAA

[471] Joshua Luczak. On how to approach the approach to equilibrium. Philosophy of Science, 83(3):393–411, July 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

MacLeod:2016:MCR

[472] Miles MacLeod and Michiru Nagatsu. Model coupling in resource economics: Conditions for effective interdisciplinary collaboration. Philosophy of Science, 83(3):412–433, July 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

deVignemont:2016:BEP

[473] Frédérique de Vignemont and Pierre Jacob. Beyond empathy for pain. Philosophy of Science, 83(3):434–445, July 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Winning:2016:RBA

[474] Jason Winning and William Bechtel. Review of Biological Autonomy. Philosophy of Science, 83(3):446–452, July 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Schindler:2016:TE

[475] Samuel Schindler. A theory of everything. Philosophy of Science, 83(3): 453–458, July 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anonymous:2016:Mb

[476] Anonymous. Masthead. Philosophy of Science, 83(4):??, October 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Sterkenburg:2016:SPO

[477] Tom F. Sterkenburg. Solomonoff prediction and Occam’s Razor. Philosophy of Science, 83(4):459–479, October 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Elliott:2016:EOA

[478] Katrina Elliott. Explaining (one aspect of) the principal principle without (much) metaphysics. Philosophy of Science, 83(4):480–499, October 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

deMelo-Martin:2016:RUI

[479] Inmaculada de Melo-Martín and Kristen Intemann. The risk of using inductive risk to challenge the value-free ideal. Philosophy of Science, 83 (4):500–520, October 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). See comment [631].

Sterelny:2016:CH

[480] Kim Sterelny. Contingency and history. Philosophy of Science, 83(4):521– 539, October 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Davies:2016:CCI

[481] Will Davies. Color constancy, illumination, and matching. Philosophy of Science, 83(4):540–562, October 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Gryb:2016:RHA

[482] Sean Gryb and Karim P. Y. Thébault. Regarding the ‘hole argument’ and the ‘problem of time’. Philosophy of Science, 83(4):563–584, October 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Fraser:2016:SSB

[483] James D. Fraser. Spontaneous symmetry breaking in finite systems. Philosophy of Science, 83(4):585–605, October 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Steel:2016:AEI

[484] Daniel Steel. Accepting an epistemically inferior alternative? A comment on Elliott and McKaughan. Philosophy of Science, 83(4):606–612, October 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

OConnor:2016:BHB

[485] Cailin O’Connor and James Owen Weatherall. Black holes, Black–Scholes, and prairie voles: an essay review of Simulation and Similarity, by Michael Weisberg. Philosophy of Science, 83(4):613–626, October 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Humphreys:2016:MMM

[486] Paul Humphreys. Models, mathematics, and measurement: a review of Reconstructing Reality by Margaret Morrison. Philosophy of Science, 83 (4):627–633, October 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Wright:2016:ROC

[487] Wayne Wright. Review of Outside Color: Perceptual Science and the Puzzle of Color in Philosophy by Mazviita Chirimuuta. Philosophy of Science, 83(4):634–639, October 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Autzen:2016:BRE

[488] Bengt Autzen. Book review: Elliott Sober. Ockham’s Razors: a User’s Manual. Philosophy of Science, 83(4):640–643, October 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Smeenk:2016:P

[489] Chris Smeenk. Preface. Philosophy of Science, 83(5):645–646, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Longino:2016:FB

[490] Helen Longino. Foregrounding the background. Philosophy of Science, 83 (5):647–661, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Sullivan:2016:CSU

[491] Jacqueline Anne Sullivan. Construct stabilization and the unity of the mind–brain sciences. Philosophy of Science, 83(5):662–673, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Weiskopf:2016:IMR

[492] Daniel A. Weiskopf. Integrative modeling and the role of neural constraints. Philosophy of Science, 83(5):674–685, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Boone:2016:MA

[493] Worth Boone and Gualtiero Piccinini. Mechanistic abstraction. Philosophy of Science, 83(5):686–697, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Craver:2016:EPN

[494] Carl F. Craver. The explanatory power of network models. Philosophy of Science, 83(5):698–709, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Muldoon:2016:NMN

[495] Sarah Feldt Muldoon and Danielle S. Bassett. Network and multilayer network approaches to understanding human brain dynamics. Philosophy of Science, 83(5):710–720, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Potochnik:2016:SEP

[496] Angela Potochnik. Scientific explanation: Putting communication first. Philosophy of Science, 83(5):721–732, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Reutlinger:2016:TMT

[497] Alexander Reutlinger. Is there a monist theory of causal and noncausal explanations? The counterfactual theory of scientific explanation. Philosophy of Science, 83(5):733–745, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Chirimuuta:2016:VPH

[498] M. Chirimuuta. Vision, perspectivism, and haptic realism. Philosophy of Science, 83(5):746–756, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Massimi:2016:TTS

[499] Michela Massimi. Three tales of scientific success. Philosophy of Science, 83(5):757–767, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Cartwright:2016:LTK

[500] Nancy Cartwright. Loose talk kills: What’s worrying about unity of method. Philosophy of Science, 83(5):768–778, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Kourany:2016:SSK

[501] Janet A. Kourany. Should some knowledge be forbidden? The case of cognitive differences research. Philosophy of Science, 83(5):779–790, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Spencer:2016:DHC

[502] Quayshawn Spencer. Do humans have continental populations? Philosophy of Science, 83(5):791–802, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Griesemer:2016:RCL

[503] James Griesemer. Reproduction in complex life cycles: Toward a developmental reaction norms perspective. Philosophy of Science, 83(5):803–815, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Godfrey-Smith:2016:CLC

[504] Peter Godfrey-Smith. Complex life cycles and the evolutionary process. Philosophy of Science, 83(5):816–827, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Herron:2016:FIC

[505] Matthew D. Herron. Fitness and individuality in complex life cycles. Philosophy of Science, 83(5):828–834, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

OMalley:2016:REM

[506] Maureen A. O’Malley. Reproduction expanded: Multigenerational and multilineal units of evolution. Philosophy of Science, 83(5):835–847, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Birch:2016:HTC

[507] Jonathan Birch. Hamilton’s two conceptions of social fitness. Philosophy of Science, 83(5):848–860, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Queller:2016:KSD

[508] David C. Queller. Kin selection and its discontents. Philosophy of Science, 83(5):861–872, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Okasha:2016:HRI

[509] Samir Okasha. On Hamilton’s rule and inclusive fitness theory with nonadditive payoffs. Philosophy of Science, 83(5):873–883, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Forber:2016:ESR

[510] Patrick Forber and Rory Smead. The evolution of spite, recognition, and morality. Philosophy of Science, 83(5):884–896, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

OConnor:2016:EGM

[511] Cailin O’Connor. The evolution of guilt: a model-based approach. Philosophy of Science, 83(5):897–908, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Downes:2016:CVS

[512] Stephen M. Downes. Confronting variation in the social and behavioral sciences. Philosophy of Science, 83(5):909–920, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Kaiser:2016:LCM

[513] Marie I. Kaiser. On the limits of causal modeling: Spatially-structurally complex biological phenomena. Philosophy of Science, 83(5):921–933, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Gebharter:2016:MAM

[514] Alexander Gebharter and Gerhard Schurz. A modeling approach for mechanisms featuring causal cycles. Philosophy of Science, 83(5):934–945, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Casini:2016:HMM

[515] Lorenzo Casini. How to model mechanistic hierarchies. Philosophy of Science, 83(5):946–958, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Weber:2016:IDB

[516] Marcel Weber. On the incompatibility of dynamical biological mechanisms and causal graphs. Philosophy of Science, 83(5):959–971, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Redei:2016:MTA

[517] Miklós Rédei and Zalán Gyenis. Measure theoretic analysis of consistency of the principal principle. Philosophy of Science, 83(5):972–987, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Bana:2016:FCP

[518] Gergei Bana. On the formal consistency of the principal principle. Philosophy of Science, 83(5):988–1001, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Castellani:2016:WCP

[519] Elena Castellani and Jenann Ismael. Which Curie’s principle? Philosophy of Science, 83(5):1002–1013, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Norton:2016:CT

[520] John D. Norton. Curie’s truism. Philosophy of Science, 83(5):1014–1026, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Fox:2016:NEP

[521] Craig W. Fox. The Newtonian equivalence principle: How the relativity of acceleration led Newton to the equivalence of inertial and gravitational mass. Philosophy of Science, 83(5):1027–1038, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Weatherall:2016:UG

[522] James Owen Weatherall. Understanding gauge. Philosophy of Science, 83(5):1039–1049, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). See erratum [643].

Manchak:2016:GIT

[523] John Byron Manchak. On Gödel and the ideality of time. Philosophy of Science, 83(5):1050–1058, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

deHaro:2016:RBD

[524] Sebastian de Haro, Nicholas Teh, and Jeremy Butterfield. On the relation between dualities and gauge symmetries. Philosophy of Science, 83(5): 1059–1069, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Hendry:2016:SA

[525] Robin Findlay Hendry. Structure as abstraction. Philosophy of Science, 83(5):1070–1081, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Fagan:2016:IOC

[526] Melinda Bonnie Fagan. Interventionist omissions: a critical case study of mechanistic explanation in biology. Philosophy of Science, 83(5):1082– 1097, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Alexandrova:2016:CVV

[527] Anna Alexandrova and Daniel M. Haybron. Is construct validation valid? Philosophy of Science, 83(5):1098–1109, December 2016. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Thompson:2016:EJC

[528] Erica Thompson, Roman Frigg, and Casey Helgeson. Expert judgment for climate change adaptation. Philosophy of Science, 83(5):1110–1121, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Goodwin:2016:MFF

[529] William M. Goodwin and Eric Winsberg. Missing the forest for the fish: How much does the ‘Hawkmoth Effect’ threaten the viability of climate projections? Philosophy of Science, 83(5):1122–1132, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Steele:2016:DMT

[530] Katie Steele and Charlotte Werndl. The diversity of model tuning practices in climate science. Philosophy of Science, 83(5):1133–1144, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anonymous:2016:Mc

[531] Anonymous. Masthead. Philosophy of Science, 83(5):??, December 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anderson:2017:SAW

[532] Wes Anderson. Some adaptations were not positive causal factors for reproductive success. Philosophy of Science, 84(1):1–13, January 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Lynch:2017:IHC

[533] Kate E. Lynch and Pierrick Bourrat. Interpreting heritability causally. Philosophy of Science, 84(1):14–34, January 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Driscoll:2017:ECC

[534] Catherine Driscoll. The evolutionary culture concepts. Philosophy of Science, 84(1):35–55, January 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Festa:2017:UGB

[535] Roberto Festa and Gustavo Cevolani. Unfolding the grammar of Bayesian confirmation: Likelihood and antilikelihood principles. Philosophy of Science, 84(1):56–81, January 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Schippers:2017:RTA

[536] Michael Schippers. A representation theorem for absolute confirmation. Philosophy of Science, 84(1):82–91, January 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Myrvold:2017:EIU

[537] Wayne C. Myrvold. On the evidential import of unification. Philosophy of Science, 84(1):92–114, January 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Bramson:2017:UPM

[538] Aaron Bramson, Patrick Grim, Daniel J. Singer, William J. Berger, Graham Sack, Steven Fisher, Carissa Flocken, and Bennett Holman. Understanding polarization: Meanings, measures, and model evaluation. Philosophy of Science, 84(1):115–159, January 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Franco:2017:ANV

[539] Paul L. Franco. Assertion, nonepistemic values, and scientific practice. Philosophy of Science, 84(1):160–180, January 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Acerbi:2017:BRT

[540] Alberto Acerbi. Book review: Tim Lewens. Cultural Evolution: Conceptual Challenges. Philosophy of Science, 84(1):181–184, January 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Wray:2017:BRR

[541] K. Brad Wray. Book review: Robert J. Richards and Lorraine Daston. Kuhn’s Structure of Scientific Revolutions at Fifty: Reflections on a Science Classic. Philosophy of Science, 84(1):184–188, January 2017. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anonymous:2017:Ma

[542] Anonymous. Masthead. Philosophy of Science, 84(1):??, January 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

DeBona:2017:GIA

[543] Glauber De Bona and Julia Staffel. Graded incoherence for accuracyfirsters. Philosophy of Science, 84(2):189–213, April 2017. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Baumgartner:2017:ATC

[544] Michael Baumgartner and Lorenzo Casini. An abductive theory of constitution. Philosophy of Science, 84(2):214–233, April 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Rosenstock:2017:ENL

[545] Sarita Rosenstock, Justin Bruner, and Cailin O’Connor. In epistemic networks, is less really more? Philosophy of Science, 84(2):234–252, April 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Bechtel:2017:ETC

[546] William Bechtel. Explicating top-down causation using networks and dynamics. Philosophy of Science, 84(2):253–274, April 2017. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Doolittle:2017:MMC

[547] W. Ford Doolittle. Making the most of clade selection. Philosophy of Science, 84(2):275–295, April 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Jackson:2017:CEP

[548] John P. Jackson, Jr. Cognitive/evolutionary psychology and the history of racism. Philosophy of Science, 84(2):296–314, April 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Roberts:2017:TMA

[549] Bryan W. Roberts. Three myths about time reversal in quantum theory. Philosophy of Science, 84(2):315–334, April 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Hancox-Li:2017:SCF

[550] Leif Hancox-Li. Solutions in constructive field theory. Philosophy of Science, 84(2):335–358, April 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Climenhaga:2017:HEG

[551] Nevin Climenhaga. How explanation guides confirmation. Philosophy of Science, 84(2):359–368, April 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Stanford:2017:BTJ

[552] P. Kyle Stanford. Bending toward justice. Philosophy of Science, 84(2): 369–376, April 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Pham:2017:RMM

[553] Michelle Pham. Review of Making Medical Knowledge. Philosophy of Science, 84(2):377–384, April 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anonymous:2017:RPS

[554] Anonymous. Referees for Philosophy of Science. Philosophy of Science, 84(2):385–390, April 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anonymous:2017:Mb

[555] Anonymous. Masthead. Philosophy of Science, 84(2):??, April 2017. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Roche:2017:EUN

[556] William Roche and Elliott Sober. Explanation = unification? A new criticism of Friedman’s theory and a reply to an old one. Philosophy of Science, 84(3):391–413, July 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Stegenga:2017:RIE

[557] Jacob Stegenga and Tarun Menon. Robustness and independent evidence. Philosophy of Science, 84(3):414–435, July 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Loew:2017:ACD

[558] Christian Loew. The asymmetry of counterfactual dependence. Philosophy of Science, 84(3):436–455, July 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Schurz:2017:ICR

[559] Gerhard Schurz. Interactive causes: Revising the Markov condition. Philosophy of Science, 84(3):456–479, July 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Emery:2017:NGO

[560] Nina Emery. A naturalist’s guide to objective chance. Philosophy of Science, 84(3):480–499, July 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). See reply [754].

Bradley:2017:CCA

[561] Richard Bradley, Casey Helgeson, and Brian Hill. Climate change assessments: Confidence, probability, and decision. Philosophy of Science, 84(3):500–522, July 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Garson:2017:GSE

[562] Justin Garson. A generalized selected effects theory of function. Philosophy of Science, 84(3):523–543, July 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Galeazzi:2017:SRR

[563] Paolo Galeazzi and Michael Franke. Smart representations: Rationality and evolution in a richer environment. Philosophy of Science, 84(3):544– 573, July 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Weber:2017:WKC

[564] Marcel Weber. Which kind of causal specificity matters biologically? Philosophy of Science, 84(3):574–585, July 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Boyd:2017:FFG

[565] Nora Mills Boyd. Franklin’s field guide to scientific experiments. Philosophy of Science, 84(3):586–594, July 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Fletcher:2017:ATE

[566] Samuel C. Fletcher. Against the topologists: Essay review of New Foundations for Physical Geometry. Philosophy of Science, 84(3):595–603, July 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Shagrir:2017:RPC

[567] Oron Shagrir. Review of Physical Computation: a Mechanistic Account by Gualtiero Piccinini. Philosophy of Science, 84(3):604–612, July 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anonymous:2017:Mc

[568] Anonymous. Masthead. Philosophy of Science, 84(3):??, July 2017. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Levinstein:2017:PGE

[569] Benjamin Anders Levinstein. A Pragmatist’s guide to epistemic utility. Philosophy of Science, 84(4):613–638, October 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Epstein:2017:FTA

[570] Peter Fisher Epstein. The fine-tuning argument and the requirement of total evidence. Philosophy of Science, 84(4):639–658, October 2017. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Howson:2017:PGS

[571] Colin Howson. Putting on the Garber style? Better not. Philosophy of Science, 84(4):659–676, October 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Belkind:2017:NI

[572] Ori Belkind. On Newtonian induction. Philosophy of Science, 84(4):677– 697, October 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Heesen:2017:CIS

[573] Remco Heesen. Communism and the incentive to share in science. Philosophy of Science, 84(4):698–716, October 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Glymour:2017:CUC

[574] Bruce Glymour. Cross-unit causation and the identity of groups. Philosophy of Science, 84(4):717–736, October 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Raz:2017:VPG

[575] Tim Räz. The Volterra Principle generalized. Philosophy of Science, 84 (4):737–760, October 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Echeverri:2017:VRI

[576] Santiago Echeverri. Visual reference and iconic content. Philosophy of Science, 84(4):761–781, October 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Currie:2017:RER

[577] Adrian Currie. Review of Evidential Reasoning in Archaeology. Philosophy of Science, 84(4):782–790, October 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Heilmann:2017:BRM

[578] Conrad Heilmann. Book review: Marcel Boumans, Science Outside the Laboratory: Measurement in Field Science and Economics. Philosophy of Science, 84(4):791–794, October 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anonymous:2017:Md

[579] Anonymous. Masthead. Philosophy of Science, 84(4):??, October 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anonymous:2017:Me

[580] Anonymous. Masthead. Philosophy of Science, 84(4):??, October 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Parker:2017:P

[581] Wendy S. Parker. Preface. Philosophy of Science, 84(5):795–796, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Romeijn:2017:ICP

[582] Jan-Willem Romeijn. Inherent complexity: a problem for statistical model evaluation. Philosophy of Science, 84(5):797–809, December 2017. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Schupbach:2017:HCB

[583] Jonah N. Schupbach and David H. Glass. Hypothesis competition beyond mutual exclusivity. Philosophy of Science, 84(5):810–824, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Schurz:2017:NFL

[584] Gerhard Schurz. No free lunch theorem, inductive skepticism, and the optimality of meta-induction. Philosophy of Science, 84(5):825–839, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Sprenger:2017:TAA

[585] Jan Sprenger and Jacob Stegenga. Three arguments for absolute outcome measures. Philosophy of Science, 84(5):840–852, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Vieland:2017:MSE

[586] Veronica J. Vieland. Measurement of statistical evidence: Picking up where Hacking and others left off. Philosophy of Science, 84(5):853–865, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Walsh:2017:NCP

[587] Kirsten Walsh. Newton: From certainty to probability? Philosophy of Science, 84(5):866–878, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Burnston:2017:RPB

[588] Daniel C. Burnston. Real patterns in biological explanation. Philosophy of Science, 84(5):879–891, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Saatsi:2017:DST

[589] Juha Saatsi. Dynamical systems theory and explanatory indispensability. Philosophy of Science, 84(5):892–904, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Skow:2017:LRC

[590] Bradford Skow. Levels of reasons and causal explanation. Philosophy of Science, 84(5):905–915, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Wakil:2017:MEB

[591] Samantha Wakil and James Justus. Mathematical explanation and the biological optimality fallacy. Philosophy of Science, 84(5):916–930, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Hicks:2017:MFF

[592] Michael Townsen Hicks. Making fit fit. Philosophy of Science, 84(5): 931–943, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Martinez:2017:SPC

[593] Eduardo J. Martinez. Stable property clusters and their grounds. Philosophy of Science, 84(5):944–955, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). See erratum [632].

Tan:2017:ICR

[594] Peter Tan. Interventions and counternomic reasoning. Philosophy of Science, 84(5):956–969, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Boesch:2017:TSP

[595] Brandon Boesch. There Is a special problem of scientific representation. Philosophy of Science, 84(5):970–981, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Nguyen:2017:SRT

[596] James Nguyen. Scientific representation and theoretical equivalence. Philosophy of Science, 84(5):982–995, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Tulodziecki:2017:ASR

[597] Dana Tulodziecki. Against selective realism(s). Philosophy of Science, 84 (5):996–1007, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Holman:2017:EIS

[598] Bennett Holman and Justin Bruner. Experimentation by industrial selection. Philosophy of Science, 84(5):1008–1019, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

LeBihan:2017:EDD

[599] Soazig Le Bihan and Iheanyi Amadi. On epistemically detrimental dissent: Contingent enabling factors versus stable difference-makers. Philosophy of Science, 84(5):1020–1030, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Romero:2017:NVR

[600] Felipe Romero. Novelty versus replicability: Virtues and vices in the reward system of science. Philosophy of Science, 84(5):1031–1043, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Schroeder:2017:UDV

[601] S. Andrew Schroeder. Using democratic values in science: an objection and (partial) response. Philosophy of Science, 84(5):1044–1054, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Bourrat:2017:DMH

[602] Pierrick Bourrat and Qiaoying Lu. Dissolving the missing heritability problem. Philosophy of Science, 84(5):1055–1067, December 2017. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Brusse:2017:RRD

[603] Carl Brusse and Justin Bruner. Responsiveness and robustness in the David Lewis signaling game. Philosophy of Science, 84(5):1068–1079, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Fulda:2017:NSM

[604] Fermín C. Fulda. Natural selection, mechanism, and the statistical interpretation. Philosophy of Science, 84(5):1080–1092, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Garson:2017:AOF

[605] Justin Garson. Against organizational functions. Philosophy of Science, 84(5):1093–1103, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

McConwell:2017:CIP

[606] Alison K. McConwell. Contingency and individuality: a plurality of evolutionary individuality types. Philosophy of Science, 84(5):1104–1116, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Molter:2017:MI

[607] Dan Molter. On mushroom individuality. Philosophy of Science, 84(5): 1117–1127, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Otsuka:2017:CHC

[608] Jun Otsuka. The causal homology concept. Philosophy of Science, 84(5): 1128–1139, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Chirimuuta:2017:CTE

[609] M. Chirimuuta. Crash testing an engineering framework in neuroscience: Does the idea of robustness break down? Philosophy of Science, 84(5): 1140–1151, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Dacey:2017:ACB

[610] Mike Dacey. Anthropomorphism as cognitive bias. Philosophy of Science, 84(5):1152–1164, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Feest:2017:POR

[611] Uljana Feest. Phenomena and objects of research in the cognitive and behavioral sciences. Philosophy of Science, 84(5):1165–1176, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Pence:2017:PCC

[612] David Evan Pence. Potential controversies: Causation and the Hodgkin and Huxley equations. Philosophy of Science, 84(5):1177–1188, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Dougherty:2017:SSG

[613] John Dougherty. Sameness and separability in gauge theories. Philosophy of Science, 84(5):1189–1201, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Lehmkuhl:2017:LVC

[614] Dennis Lehmkuhl. Literal versus careful interpretations of scientific theories: The vacuum approach to the problem of motion in general relativity. Philosophy of Science, 84(5):1202–1214, December 2017. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Manchak:2017:IST

[615] John Byron Manchak. On the inextendibility of space–time. Philosophy of Science, 84(5):1215–1225, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Martens:2017:RCA

[616] Niels C. M. Martens. Regularity comparativism about mass in Newtonian gravity. Philosophy of Science, 84(5):1226–1238, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

McCoy:2017:CTA

[617] C. D. McCoy. Can typicality arguments dissolve Cosmology’s flatness problem? Philosophy of Science, 84(5):1239–1252, December 2017. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Moller-Nielsen:2017:IIM

[618] Thomas Møller-Nielsen. Invariance, interpretation, and motivation. Philosophy of Science, 84(5):1253–1264, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Roberts:2017:UO

[619] Bryan W. Roberts. Unreal observables. Philosophy of Science, 84(5): 1265–1274, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

deOlano:2017:ICS

[620] Pablo Ruiz de Olano. Intimate connections: Symmetries and conservation laws in quantum versus classical mechanics. Philosophy of Science, 84(5): 1275–1288, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Werndl:2017:MGB

[621] Charlotte Werndl and Roman Frigg. Mind the gap: Boltzmannian versus Gibbsian equilibrium. Philosophy of Science, 84(5):1289–1302, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Zwier:2017:ICT

[622] Karen R. Zwier. Interventionist causation in thermodynamics. Philosophy of Science, 84(5):1303–1315, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anonymous:2017:Mf

[623] Anonymous. Masthead. Philosophy of Science, 84(5):??, December 2017. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Rubin:2018:DIF

[624] Hannah Rubin. The debate over inclusive fitness as a debate over methodologies. Philosophy of Science, 85(1):1–30, January 2018. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Ludwig:2018:LGN

[625] David Ludwig. Letting go of “natural kind”: Toward a multidimensional framework of nonarbitrary classification. Philosophy of Science, 85(1):31– 52, January 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Cisewski:2018:SMB

[626] Jessi Cisewski, Joseph B. Kadane, Mark J. Schervish, Teddy Seidenfeld, and Rafael Stern. Standards for modest Bayesian credences. Philosophy of Science, 85(1):53–78, January 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Icard:2018:BBR

[627] Thomas F. Icard. Bayes, bounds, and rational analysis. Philosophy of Science, 85(1):79–101, January 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Feintzeig:2018:CAQ

[628] Benjamin H. Feintzeig. On the choice of algebra for quantization. Philosophy of Science, 85(1):102–125, January 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Kao:2018:OED

[629] Molly Kao. Old evidence in the development of quantum theory. Philosophy of Science, 85(1):126–143, January 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Zahle:2018:VDC

[630] Julie Zahle. Values and data collection in social research. Philosophy of Science, 85(1):144–163, January 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Hicks:2018:IRR

[631] Daniel J. Hicks. Inductive risk and regulatory toxicology: a comment on de Melo–Martín and Intemann. Philosophy of Science, 85(1):164–174, January 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). See [479].

Anonymous:2018:EEJ

[632] Anonymous. Erratum: Eduardo J. Martinez’s “Stable Property Clusters and Their Grounds” (in vol. 84, no. 5, December 2017). Philosophy of Science, 85(1):??, January 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). See [593].

Anonymous:2018:Ma

[633] Anonymous. Masthead. Philosophy of Science, 85(1):??, January 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Green:2018:WDO

[634] E. J. Green. What do object files pick out? Philosophy of Science, 85(2): 177–200, April 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Teh:2018:RRR

[635] Nicholas J. Teh. Recovering recovery: On the relationship between gauge symmetry and Trautman recovery. Philosophy of Science, 85(2):201–224, April 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Franklin:2018:RGE

[636] Alexander Franklin. On the renormalization group explanation of universality. Philosophy of Science, 85(2):225–248, April 2018. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Dewar:2018:MG

[637] Neil Dewar. Maxwell gravitation. Philosophy of Science, 85(2):249–270, April 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Jhun:2018:WPC

[638] Jennifer S. Jhun. What’s the point of Ceteris Paribus? or, how to understand supply and demand curves. Philosophy of Science, 85(2):271–292, April 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Pomatto:2018:ATI

[639] Luciano Pomatto and Alvaro Sandroni. An axiomatic theory of inductive inference. Philosophy of Science, 85(2):293–315, April 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Marxen:2018:BRR

[640] Chad Marxen and Gerard Rothfus. Book review: Richard Pettigrew, Accuracy and the Laws of Credence. Oxford: Oxford University Press (2016), 256 pp., $74.00 (cloth). Philosophy of Science, 85(2):316–320, April 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Tabery:2018:BRK

[641] James Tabery. Book review: Kenneth F. Schaffner, Behaving: What’s Genetic, What’s Not, and Why Should We Care? New York: Oxford University Press (2016), xiii + 287 pp., $78.00 (cloth). Philosophy of Science, 85(2):321–324, April 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anonymous:2018:RPS

[642] Anonymous. Referees for Philosophy of Science. Philosophy of Science, 85(2):326–333, April 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anonymous:2018:EJO

[643] Anonymous. Erratum: James Owen Weatherall’s Understanding Gauge (Philosophy of Science 83, no. 5 [2016]: 1039–49). Philosophy of Science, 85(2):??, April 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). See [522].

Anonymous:2018:Mb

[644] Anonymous. Masthead. Philosophy of Science, 85(2):??, April 2018. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anonymous:2018:PM

[645] Anonymous. Perspectival modeling. Philosophy of Science, 85(3):335– 359, July 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Biddle:2018:AZV

[646] Justin B. Biddle. “Antiscience zealotry”? Values, epistemic risk, and the GMO debate. Philosophy of Science, 85(3):360–379, July 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Rubin:2018:DCS

[647] Hannah Rubin and Cailin O’Connor. Discrimination and collaboration in science. Philosophy of Science, 85(3):380–402, July 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Boyd:2018:EE

[648] Nora Mills Boyd. Evidence enriched. Philosophy of Science, 85(3):403– 421, July 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Stroing:2018:DEE

[649] Pascal Ströing. Data, evidence, and explanatory power. Philosophy of Science, 85(3):422–441, July 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Cohen:2018:EJC

[650] Michael Cohen. Explanatory justice: The case of disjunctive explanations. Philosophy of Science, 85(3):442–454, July 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Saatsi:2018:TRL

[651] Juha Saatsi and Alexander Reutlinger. Taking reductionism to the limit: How to rebut the antireductionist argument from infinite limits. Philosophy of Science, 85(3):455–482, July 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Loew:2018:FTA

[652] Christian Loew. Fundamentality and time’s arrow. Philosophy of Science, 85(3):483–500, July 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Lean:2018:ISE

[653] Christopher Hunter Lean. Indexically structured ecological communities. Philosophy of Science, 85(3):501–522, July 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anonymous:2018:NPR

[654] Anonymous. New perspectives on reductionism in biology. Philosophy of Science, 85(3):523–529, July 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Swanson:2018:RJB

[655] Noel Swanson. Review of Jonathan Bain’s CPT Invariance and the Spin-Statistics Connection. Philosophy of Science, 85(3):530–539, July 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anonymous:2018:RSL

[656] Anonymous. Review of Sabina Leonelli’s Data-Centric Biology: a Philosophical Study. Philosophy of Science, 85(3):540–550, July 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anonymous:2018:Mc

[657] Anonymous. Masthead. Philosophy of Science, 85(3):??, July 2018. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Ross:2018:CSP

[658] Lauren N. Ross. Causal selection and the pathway concept. Philosophy of Science, 85(4):551–572, October 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Bechtel:2018:ICC

[659] William Bechtel. The importance of constraints and control in biological mechanisms: Insights from cancer research. Philosophy of Science, 85 (4):573–593, October 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Bausman:2018:MNN

[660] William C. Bausman. Modeling: Neutral, null, and baseline. Philosophy of Science, 85(4):594–616, October 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Barrett:2018:WDS

[661] Thomas William Barrett. What do symmetries tell us about structure? Philosophy of Science, 85(4):617–639, October 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Pober:2018:WER

[662] Jeremy Pober. What emotions really are (in the theory of constructed emotions). Philosophy of Science, 85(4):640–659, October 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Angner:2018:WPR

[663] Erik Angner. What preferences really are. Philosophy of Science, 85 (4):660–681, October 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Blanchard:2018:BEU

[664] Thomas Blanchard. Bayesianism and explanatory unification: a compatibilist account. Philosophy of Science, 85(4):682–703, October 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

ChoGlueck:2018:EGS

[665] Christopher ChoGlueck. The error is in the gap: Synthesizing accounts for societal values in science. Philosophy of Science, 85(4):704–725, October 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Garson:2018:RKN

[666] Justin Garson. Review of Karen Neander’s A Mark of the Mental: In Defense of Informational Teleosemantics. Philosophy of Science, 85(4): 726–734, October 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Allori:2018:PJL

[667] Valia Allori. Book review: Peter J. Lewis, Quantum Ontology: A Guide to the Metaphysics of Quantum Mechanics. Oxford: Oxford University Press (2016), 207 pp., $35.00 (paper). Philosophy of Science, 85(4):735–738, October 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anonymous:2018:Md

[668] Anonymous. Masthead. Philosophy of Science, 85(4):??, October 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Parker:2018:P

[669] Wendy S. Parker. Preface. Philosophy of Science, 85(5):739–740, December 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Leonelli:2018:TDT

[670] Sabina Leonelli. The time of data: Timescales of data use in the life sciences. Philosophy of Science, 85(5):741–754, December 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

McAllister:2018:SRO

[671] James W. McAllister. Scientists’ reuse of old empirical data: Epistemological aspects. Philosophy of Science, 85(5):755–766, December 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Fuller:2018:MRE

[672] Jonathan Fuller. Meta-research evidence for evaluating therapies. Philosophy of Science, 85(5):767–780, December 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Keren:2018:PUW

[673] Arnon Keren. The public understanding of what? Laypersons’ epistemic needs, the division of cognitive labor, and the demarcation of science. Philosophy of Science, 85(5):781–792, December 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Bokulich:2018:REE

[674] Alisa Bokulich. Representing and explaining: The eikonic conception of scientific explanation. Philosophy of Science, 85(5):793–805, December 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Pincock:2018:ERC

[675] Christopher Pincock. Explanatory relevance and contrastive explanation. Philosophy of Science, 85(5):806–818, December 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anderl:2018:SSA

[676] Sibylle Anderl. Simplicity and simplification in astrophysical modeling. Philosophy of Science, 85(5):819–831, December 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Lenhard:2018:HEM

[677] Johannes Lenhard. Holism, or the erosion of modularity: a methodological challenge for validation. Philosophy of Science, 85(5):832–844, December 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Ereshefsky:2018:NKM

[678] Marc Ereshefsky. Natural kinds, mind independence, and defeasibility. Philosophy of Science, 85(5):845–856, December 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Elliott:2018:AIF

[679] Kevin C. Elliott. Addressing industry-funded research with criteria for objectivity. Philosophy of Science, 85(5):857–868, December 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Holman:2018:SDC

[680] Bennett Holman and Sally Geislar. Sex drugs and corporate ventriloquism: How to evaluate science policies intended to manage industry-funded bias. Philosophy of Science, 85(5):869–881, December 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Pinto:2018:DSI

[681] Manuela Fernández Pinto. Democratizing strategies for industry-funded medical research: a cautionary tale. Philosophy of Science, 85(5):882–894, December 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Steel:2018:WTV

[682] Daniel Steel. Wishful thinking and values in science. Philosophy of Science, 85(5):895–905, December 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Palmerino:2018:DWW

[683] Carla Rita Palmerino. Discussing what would happen: The role of thought experiments in Galileo’s dialogues. Philosophy of Science, 85(5):906–918, December 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

VanDyck:2018:IGP

[684] Maarten Van Dyck. Idealization and Galileo’s proto-inertial principle. Philosophy of Science, 85(5):919–929, December 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Pence:2018:HDD

[685] Charles H. Pence and Grant Ramsey. How to do digital philosophy of science. Philosophy of Science, 85(5):930–941, December 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Plutynski:2018:WHD

[686] Anya Plutynski and Marta Bertolaso. What and how do cancer systems biologists explain? Philosophy of Science, 85(5):942–954, December 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Jones:2018:SEA

[687] Nicholaos Jones. Strategies of explanatory abstraction in molecular systems biology. Philosophy of Science, 85(5):955–968, December 2018. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Huneman:2018:NST

[688] Philippe Huneman. Neutral spaces and topological explanations in evolutionary biology: Lessons from some landscapes and mappings. Philosophy of Science, 85(5):969–983, December 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Brigandt:2018:EMP

[689] Ingo Brigandt. Explanation of molecular processes without tracking mechanism operation. Philosophy of Science, 85(5):984–997, December 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Green:2018:SDD

[690] Sara Green. Scale dependency and downward causation in biology. Philosophy of Science, 85(5):998–1011, December 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Pocheville:2018:BIC

[691] Arnaud Pocheville. Biological information as choice and construction. Philosophy of Science, 85(5):1012–1025, December 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Birch:2018:ACH

[692] Jonathan Birch. Animal cognition and human values. Philosophy of Science, 85(5):1026–1037, December 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Scarantino:2018:EES

[693] Andrea Scarantino. Emotional expressions as speech act analogs. Philosophy of Science, 85(5):1038–1053, December 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Piccinini:2018:EPA

[694] Gualtiero Piccinini and Armin Schulz. The evolution of psychological altruism. Philosophy of Science, 85(5):1054–1064, December 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Bickle:2018:MOI

[695] John Bickle. From microscopes to optogenetics: Ian Hacking vindicated. Philosophy of Science, 85(5):1065–1077, December 2018. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Robins:2018:MOI

[696] Sarah K. Robins. Memory and optogenetic intervention: Separating the engram from the ecphory. Philosophy of Science, 85(5):1078–1089, December 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Sullivan:2018:OPP

[697] Jacqueline Anne Sullivan. Optogenetics, pluralism, and progress. Philosophy of Science, 85(5):1090–1101, December 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Elliott-Graves:2018:GCI

[698] Alkistis Elliott-Graves. Generality and causal interdependence in ecology. Philosophy of Science, 85(5):1102–1114, December 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Eliot:2018:EIC

[699] Christopher H. Eliot. Ecological interdependence via constraints. Philosophy of Science, 85(5):1115–1126, December 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Millstein:2018:ULC

[700] Roberta L. Millstein. Understanding Leopold’s concept of “interdependence” for environmental ethics and conservation biology. Philosophy of Science, 85(5):1127–1139, December 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Northcott:2018:EQE

[701] Robert Northcott. The efficiency question in economics. Philosophy of Science, 85(5):1140–1151, December 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Weatherall:2018:PLB

[702] James Owen Weatherall. The peculiar logic of the Black–Scholes model. Philosophy of Science, 85(5):1152–1163, December 2018. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Fraser:2018:RFS

[703] James D. Fraser. Renormalization and the formulation of scientific realism. Philosophy of Science, 85(5):1164–1175, December 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Ruetsche:2018:RGR

[704] Laura Ruetsche. Renormalization group realism: The ascent of pessimism. Philosophy of Science, 85(5):1176–1189, December 2018. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Huggett:2018:TES

[705] Nick Huggett and Christian Wüthrich. The (a)temporal emergence of spacetime. Philosophy of Science, 85(5):1190–1203, December 2018. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Weisberg:2018:DGH

[706] Michael Weisberg, Melissa Jacquart, Barry Madore, and Marja Seidel. The dark galaxy hypothesis. Philosophy of Science, 85(5):1204–1215, December 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anonymous:2018:Me

[707] Anonymous. Masthead. Philosophy of Science, 85(5):??, December 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Martens:2019:IFM

[708] Johannes Martens. Inclusive fitness as a measure of biological utility. Philosophy of Science, 86(1):1–22, January 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Stanley:2019:NMF

[709] Matthew L. Stanley, Bryce Gessell, and Felipe De Brigard. Network modularity as a foundation for neural reuse. Philosophy of Science, 86(1): 23–46, January 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Hudetz:2019:DCE

[710] Laurenz Hudetz. Definable categorical equivalence. Philosophy of Science, 86(1):47–75, January 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Clarke:2019:CAR

[711] Christopher Clarke. The correlation argument for reductionism. Philosophy of Science, 86(1):76–97, January 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Grim:2019:DAE

[712] Patrick Grim, Daniel J. Singer, Aaron Bramson, Bennett Holman, Sean McGeehan, and William J. Berger. Diversity, ability, and expertise in epistemic communities. Philosophy of Science, 86(1):98–123, January 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Boyer-Kassem:2019:SER

[713] Thomas Boyer-Kassem. Scientific expertise and risk aggregation. Philosophy of Science, 86(1):124–144, January 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Kinney:2019:EDP

[714] David Kinney. On the explanatory depth and pragmatic value of coarsegrained, probabilistic, causal explanations. Philosophy of Science, 86(1): 145–167, January 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Lawler:2019:LRW

[715] Insa Lawler. Levels of reasons why and answers to why questions. Philosophy of Science, 86(1):168–177, January 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Singer:2019:DRT

[716] Daniel J. Singer. Diversity, not randomness, trumps ability. Philosophy of Science, 86(1):178–191, January 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Heesen:2019:BRT

[717] Remco Heesen. Book review: Thomas Boyer-Kassem, Conor Mayo-Wilson, and Michael Weisberg, eds. Scientific Collaboration and Collective Knowledge: New Essays. Philosophy of Science, 86(1):192–198, January 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Bickle:2019:BRC

[718] John Bickle. Book review: Carl Gillett. Reduction and Emergence in Science and Philosophy. Philosophy of Science, 86(1):198–201, January 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anonymous:2019:Ma

[719] Anonymous. Masthead. Philosophy of Science, 86(1):??, January 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Schickore:2019:SFE

[720] Jutta Schickore. The structure and function of experimental control in the life sciences. Philosophy of Science, 86(2):203–218, April 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

vanEersel:2019:EER

[721] Gerdien G. van Eersel, Gabriela V. Koppenol-Gonzalez, and Julian Reiss. Extrapolation of experimental results through analogical reasoning from latent classes. Philosophy of Science, 86(2):219–235, April 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Stewart:2019:AAC

[722] Rush T. Stewart and Michael Nielsen. Another approach to consensus and maximally informed opinions with increasing evidence. Philosophy of Science, 86(2):236–254, April 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Kaivanto:2019:AIT

[723] Kim Kaivanto and Daniel Steel. Adjusting inferential thresholds to reflect nonepistemic values. Philosophy of Science, 86(2):255–285, April 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

McCutcheon:2019:FLS

[724] Randall G. McCutcheon. In favor of logarithmic scoring. Philosophy of Science, 86(2):286–303, April 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Gasparri:2019:PSO

[725] Luca Gasparri. Phonetic segments and the organization of speech. Philosophy of Science, 86(2):304–324, April 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

OLoughlin:2019:ATS

[726] Ian O’Loughlin and Kate McCallum. The aesthetics of theory selection and the logics of art. Philosophy of Science, 86(2):325–343, April 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Suarez:2019:RSC

[727] Mauricio Suárez and Francesca Pero. The representational semantic conception. Philosophy of Science, 86(2):344–365, April 2019. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Howson:2019:BWF

[728] Colin Howson. A better way of framing Williamson’s coin-tossing argument, but it still does not work. Philosophy of Science, 86(2):366–374, April 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Valde:2019:BRD

[729] Katherine Valde. Book review: Daniel J. Nicholson and John Dupré, eds. Everything Flows: Towards a Processual Philosophy of Biology. Philosophy of Science, 86(2):375–378, April 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Wieten:2019:KET

[730] Sarah Wieten. Book review: Kevin Elliott. A Tapestry of Values: an Introduction to Values in Science. Philosophy of Science, 86(2):378–383, April 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anonymous:2019:RPS

[731] Anonymous. Referees for Philosophy of Science. Philosophy of Science, 86(2):384–390, April 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anonymous:2019:Mb

[732] Anonymous. Masthead. Philosophy of Science, 86(2):??, April 2019. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Wilson:2019:IIA

[733] Robert A. Wilson. Incest, incest avoidance, and attachment: Revisiting the Westermarck effect. Philosophy of Science, 86(3):391–411, July 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Witteveen:2019:NSC

[734] Joeri Witteveen. Natural selection and contrastive explanation. Philosophy of Science, 86(3):412–430, July 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Mollo:2019:TTF

[735] Dimitri Coelho Mollo. Are there teleological functions to compute? Philosophy of Science, 86(3):431–452, July 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Guttinger:2019:NAR

[736] Stephan Guttinger. A new account of replication in the experimental life sciences. Philosophy of Science, 86(3):453–471, July 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Francois:2019:AVS

[737] Jordan François. Artificial versus substantial gauge symmetries: a criterion and an application to the electroweak model. Philosophy of Science, 86(3):472–496, July 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Mohseni:2019:SSD

[738] Aydin Mohseni. Stochastic stability and disagreements between dynamics. Philosophy of Science, 86(3):497–521, July 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Babic:2019:TER

[739] Boris Babic. A theory of epistemic risk. Philosophy of Science, 86(3):522– 550, July 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Robinson:2019:EPS

[740] Brian Robinson, Chad Gonnerman, and Michael O’Rourke. Experimental philosophy of science and philosophical differences across the sciences. Philosophy of Science, 86(3):551–576, July 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Burnston:2019:RAP

[741] Daniel C. Burnston. Review of Angela Potochnik’s Idealization and the Aims of Science. Philosophy of Science, 86(3):577–583, July 2019. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anonymous:2019:Mc

[742] Anonymous. Masthead. Philosophy of Science, 86(3):??, July 2019. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Waters:2019:PAP

[743] C. Kenneth Waters. Presidential Address, PSA 2016: an epistemology of scientific practice. Philosophy of Science, 86(4):585–611, October 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Palacios:2019:PTC

[744] Patricia Palacios. Phase transitions: a challenge for intertheoretic reduction? Philosophy of Science, 86(4):612–640, October 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Knuuttila:2019:DNE

[745] Tarja Knuuttila and Mary S. Morgan. Deidealization: No easy reversals. Philosophy of Science, 86(4):641–661, October 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

McLoone:2019:TIR

[746] Brian McLoone. Thumper the infinitesimal rabbit: a fictionalist perspective on some “unimaginable” model systems in biology. Philosophy of Science, 86(4):662–671, October 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Peden:2019:DIM

[747] William Peden. Direct inference in the material theory of induction. Philosophy of Science, 86(4):672–695, October 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Vassend:2019:NSB

[748] Olav Benjamin Vassend. New semantics for Bayesian inference: The interpretive problem and its solutions. Philosophy of Science, 86(4):696–718, October 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Sartenaer:2019:HBS

[749] Olivier Sartenaer. Humeanism, best system laws, and emergence. Philosophy of Science, 86(4):719–738, October 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Shan:2019:NFA

[750] Yafeng Shan. A new functional approach to scientific progress. Philosophy of Science, 86(4):739–758, October 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Woodward:2019:WSL

[751] James Woodward. On Wolfgang Spohn’s Laws of Belief. Philosophy of Science, 86(4):759–772, October 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). See [139, 752].

Spohn:2019:RJW

[752] Wolfgang Spohn. Reply to Jim Woodward’s comments on Wolfgang Spohn’s Laws of Belief. Philosophy of Science, 86(4):773–784, October 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). See [139, 751].

Lazarovici:2019:BVG

[753] Dustin Lazarovici. On Boltzmann versus Gibbs and the equilibrium in statistical mechanics. Philosophy of Science, 86(4):785–793, October 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Mitchell:2019:RNE

[754] Sam Mitchell. A reply to Nina Emery. Philosophy of Science, 86(4):794– 806, October 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). See [560].

Dorato:2019:RSP

[755] Mauro Dorato. Review of Simon Prosser’s Experiencing Time. Philosophy of Science, 86(4):807–813, October 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

dePaz:2019:BRD

[756] María de Paz. Book review: Dominique Raynaud, Scientific Controversies: A Socio-Historical Perspective on the Advancement of Science, trans. Lisa C. Chien. New Brunswick, NJ: Transaction (2015), xxii + 298 pp., $110.00 (cloth). Philosophy of Science, 86(4):814–817, October 2019. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anonymous:2019:Md

[757] Anonymous. Masthead. Philosophy of Science, 86(4):??, October 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anonymous:2019:Me

[758] Anonymous. Masthead. Philosophy of Science, 86(5):??, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Elliott:2019:P

[759] Kevin C. Elliott. Preface. Philosophy of Science, 86(5):819–820, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Eva:2019:SCS

[760] Benjamin Eva, Reuben Stern, and Stephan Hartmann. The similarity of causal structure. Philosophy of Science, 86(5):821–835, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Neal:2019:WCS

[761] Jacob P. Neal. When causal specificity does not matter (much): Insights from HIV treatment. Philosophy of Science, 86(5):836–846, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Fazekas:2019:HML

[762] Peter Fazekas and Gergely Kertesz. Are higher mechanistic levels causally autonomous? Philosophy of Science, 86(5):847–857, December 2019. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Swaim:2019:RPM

[763] Daniel G. Swaim. The roles of possibility and mechanism in narrative explanation. Philosophy of Science, 86(5):858–868, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Massimi:2019:TKE

[764] Michela Massimi. Two kinds of exploratory models. Philosophy of Science, 86(5):869–881, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Fang:2019:MEM

[765] Wei Fang. Mixed-effects modeling and nonreductive explanation. Philosophy of Science, 86(5):882–894, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Feest:2019:WRO

[766] Uljana Feest. Why replication is overrated. Philosophy of Science, 86 (5):895–905, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Frost-Arnold:2019:HHM

[767] Greg Frost-Arnold. How to be a historically motivated antirealist: The problem of misleading evidence. Philosophy of Science, 86(5):906–917, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Shech:2019:HIM

[768] Elay Shech. Historical inductions meet the material theory. Philosophy of Science, 86(5):918–929, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Isaac:2019:ELM

[769] Alistair M. C. Isaac. Epistemic loops and measurement realism. Philosophy of Science, 86(5):930–941, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Dewar:2019:SRT

[770] Neil Dewar. Supervenience, reduction, and translation. Philosophy of Science, 86(5):942–954, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Fumagalli:2019:FUE

[771] Roberto Fumagalli. (F)utility exposed. Philosophy of Science, 86(5):955– 966, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Lin:2019:HPT

[772] Hanti Lin. The hard problem of theory choice: a case study on causal inference and its faithfulness assumption. Philosophy of Science, 86(5): 967–980, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Sterkenburg:2019:MJI

[773] Tom F. Sterkenburg. The metainductive justification of induction: The pool of strategies. Philosophy of Science, 86(5):981–992, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Andreas:2019:EC

[774] Holger Andreas. Explanatory conditionals. Philosophy of Science, 86(5): 993–1004, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Franco:2019:SAT

[775] Paul L. Franco. Speech act theory and the multiple aims of science. Philosophy of Science, 86(5):1005–1015, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

DiMarco:2019:ITV

[776] Marina DiMarco and Kareem Khalifa. Inquiry tickets: Values, pursuit, and underdetermination. Philosophy of Science, 86(5):1016–1028, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Dang:2019:DCS

[777] Haixin Dang. Do collaborators in science need to agree? Philosophy of Science, 86(5):1029–1040, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

McCusker:2019:WHG

[778] Darcy McCusker. What is the harm in gendered citation practices? Philosophy of Science, 86(5):1041–1051, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Koskinen:2019:SIM

[779] Inkeri Koskinen and Kristina Rolin. Scientific/intellectual movements remedying epistemic injustice: The case of indigenous studies. Philosophy of Science, 86(5):1052–1063, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Bueter:2019:EIP

[780] Anke Bueter. Epistemic injustice and psychiatric classification. Philosophy of Science, 86(5):1064–1074, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Otsuka:2019:SM

[781] Jun Otsuka. Species as models. Philosophy of Science, 86(5):1075–1086, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Sterner:2019:ESL

[782] Beckett Sterner. Evolutionary species in light of population genomics. Philosophy of Science, 86(5):1087–1098, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Neto:2019:WL

[783] Celso Neto. What is a lineage? Philosophy of Science, 86(5):1099–1110, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Sencan:2019:TTI

[784] Sinan Sencan. A tale of two individuality accounts and integrative pluralism. Philosophy of Science, 86(5):1111–1122, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Koskinen:2019:MRB

[785] Rami Koskinen. Multiple realizability and biological modality. Philosophy of Science, 86(5):1123–1133, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Gross:2019:ORM

[786] Fridolin Gross. Occam’s Razor in molecular and systems biology. Philosophy of Science, 86(5):1134–1145, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Garson:2019:TNA

[787] Justin Garson. There are no ahistorical theories of function. Philosophy of Science, 86(5):1146–1156, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Autzen:2019:SRF

[788] Bengt Autzen. Survival, reproduction, and functional efficiency. Philosophy of Science, 86(5):1157–1167, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Powell:2019:R

[789] Russell Powell and Eric Scarffe. Rehabilitating “disease”: Function, value, and objectivity in medicine. Philosophy of Science, 86(5):1168–1178, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Klein:2019:WJO

[790] Alexander Klein. William James’s objection to epiphenomenalism. Philosophy of Science, 86(5):1179–1190, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Zerilli:2019:NRR

[791] John Zerilli. Neural redundancy and its relation to neural reuse. Philosophy of Science, 86(5):1191–1201, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Springle:2019:PRR

[792] Alison Ann Springle. Perception, representation, realism, and function. Philosophy of Science, 86(5):1202–1213, December 2019. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Martinez:2019:RRD

[793] Manolo Martínez. Representations are rate-distortion sweet spots. Philosophy of Science, 86(5):1214–1226, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Kurth:2019:EPC

[794] Charlie Kurth. Are emotions psychological constructions? Philosophy of Science, 86(5):1227–1238, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Michel:2019:MCP

[795] Matthias Michel. The mismeasure of consciousness: a problem of coordination for the perceptual awareness scale. Philosophy of Science, 86(5): 1239–1249, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Stone:2019:DDC

[796] Caroline Stone. A defense and definition of construct validity in psychology. Philosophy of Science, 86(5):1250–1261, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

McCoy:2019:DUC

[797] C. D. McCoy. Did the universe have a chance? Philosophy of Science, 86(5):1262–1272, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Menon:2019:AFD

[798] Tushar Menon. Algebraic fields and the dynamical approach to physical geometry. Philosophy of Science, 86(5):1273–1283, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Hicks:2019:WES

[799] Michael Townsen Hicks. What everyone should say about symmetries (and how Humeans get to say it). Philosophy of Science, 86(5):1284–1294, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Franklin:2019:UR

[800] Alexander Franklin. Universality reduced. Philosophy of Science, 86(5): 1295–1306, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Esser:2019:QTA

[801] Stephen Esser. The quantum theory of atoms in molecules and the interactive conception of chemical bonding. Philosophy of Science, 86(5): 1307–1317, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Pulkkinen:2019:VCH

[802] Karoliina Pulkkinen. The value of completeness: How Mendeleev used his periodic system to make predictions. Philosophy of Science, 86(5): 1318–1329, December 2019. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Broadbent:2013:PE

[803] Alex Broadbent. Philosophy of Epidemiology. New directions in the philosophy of science. Palgrave Macmillan, Basingstoke, UK, 2013. ISBN 0-230-35512-9 (hardcover). xxii + 203 pp. LCCN RA652 .B76 2013.

Craver:2013:SMD

[804] Carl F. Craver and Lindley Darden. In Search of Mechanisms: Discoveries Across the Life Sciences. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-03965-X (hardcover), 0-226-03979-X (paperback). xxii + 228 pp. LCCN QH331 .C898 2013.

Dick:2013:DCA

[805] Steven J. Dick. Discovery and Classification in Astronomy: Controversy and Consensus. Cambridge University Press, Cambridge, UK, 2013. ISBN 1-107-03361-6 (hardcover). xvi + 458 pp. LCCN QB61 .D53 2013.

Duncan:2012:CFQ

[806] Anthony Duncan. The Conceptual Framework of Quantum Field Theory. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-957326-3. x + 782 pp. LCCN QC174.45

.D86 2012. URL http://www.loc.gov/catdir/enhancements/fy1306/
2012472641-b.html;   http://www.loc.gov/catdir/enhancements/
fy1306/2012472641-d.html;   http://www.loc.gov/catdir/enhancements/
fy1306/2012472641-t.html.

Franklin:2013:SSE

[807] Allan Franklin. Shifting standards: experiments in particle physics in the twentieth century. University of Pittsburgh Press, Pittsburgh, PA, USA, 2013. ISBN 0-8229-4430-8 (hardcover). ???? pp. LCCN QC793.412 .F73 2013.

Friedman:2013:KCN

[808] Michael Friedman. Kant’s construction of nature: a reading of the metaphysical foundations of natural science. Cambridge University Press, Cambridge, UK, 2013. ISBN 0-521-19839-9 (hardcover). xix + 624 pp. LCCN B2786.Z7 F75 2013.

Godfrey-Smith:2014:PB

[809] Peter Godfrey-Smith. Philosophy of biology. Princeton foundations of contemporary philosophy. Princeton University Press, Princeton, NJ, USA, 2014. ISBN 0-691-14001-4 (hardcover). 187 pp. LCCN QH331 .G615 2014.

Hudson:2014:STP

[810] Robert Hudson. Seeing Things: the Philosophy of Reliable Observation. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-930328-2. xxiv + 274 pp. LCCN Q175.32.O27.

Kutach:2013:CBF

[811] Douglas Kutach. Causation and Its Basis in Fundamental Physics. Oxford studies in philosophy of science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-993620-X (hardcover), 0-19-993621-8 (updf). xiii + 334 pp. LCCN QC6.4.C3 K88 2013.

Pettigrew:2016:ALC

[812] Richard Pettigrew. Accuracy and the Laws of Credence. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2016. ISBN 0-19-873271-6. ix + 238 pp. LCCN BD215 .P48 2016.

Schaffner:2016:BWGa

[813] Kenneth F. Schaffner. Behaving: What’s Genetic, What’s Not, and Why Should We Care? Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2016. ISBN 0-19-517140-3 (hardcover). xiii + 287 pp. LCCN QH457 .S33 2016.

Steel:2015:PPP

[814] Daniel Steel. Philosophy and the Precautionary Principle: Science, Evidence, and Environmental Policy. Cambridge University Press, Cambridge, UK, 2015. ISBN 1-107-07816-4, 1-316-05743-7 (e-book). xv + 256 pp. LCCN Q175 .S7356 2015. URL http://www.loc.gov/catdir/

enhancements/fy1508/2015288007-b.html;   http://www.loc.gov/
catdir/enhancements/fy1508/2015288007-d.html;   http://www.loc.
gov/catdir/enhancements/fy1508/2015288007-t.html.

Strevens:2013:TIP

[815] Michael Strevens. Tychomancy: Inferring Probability from Causal Structure. Harvard University Press, Cambridge, MA, USA, 2013. ISBN 0-674-07311-8. xii + 265 pp. LCCN QC174.85.P76 S74 2013.